### v1

# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

Abstract.
We propose one finite element method for both second order linear uniformly elliptic PDE in non-divergence form and the elliptic Hamilton-Jacobi-Bellman (HJB) equation. For the linear elliptic PDE in non-divergence form, we consider two scenarios of the matrix coefficient matrix $A$. One is $A$ is uniformly continuous. The other is $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$ where $\gamma$ is a positive weight function.
We prove that optimal convergence in discrete $W^{2,p}$-norm of the numerical approximation to the strong solution for $1<p\leq 2$ on convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). If the domain is a two dimensional non-convex polygon, $p$ is valid in a neighbourhood of $\frac{4}{3}$. We also prove the well-posedness of strong solution in $W^{2,p}(\Omega)$ for both linear elliptic PDE in non-divergence form and the HJB equation for $1<p\leq 2$ on convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and for $p$ in an open interval starting from $1$ and including $\frac{4}{3}$ on two dimensional non-convex polygon. Furthermore, we relax the assumptions on the continuity of coefficients of the HJB equation, which have been widely used in literature.
Key words and phrases:
Non-divergence form, Hamilton-Jacobi-Bellman, $W^{2,p}$ estimate, Lipschitz polyhedra
2010 Mathematics Subject Classification:
65L60, 65N30, 46E35, 52B10, 26A16
1. Introduction
We mainly consider optimal convergence to the strong solution in $W_{h}^{2,p}(\Omega)$-norm (see (7)) of a finite element method for both second order linear uniformly elliptic PDE in non-divergence form (2) and the elliptic Hamilton-Jacobi-Bellman equation (3) on Lipschitz polyhedral domains in $\mathbb{R}^{d}$ ($d=2,3$). We consider the well-posedness of the strong solution in $W^{2,p}(\Omega)$ of (2) and (3) as well. The ranges of $p$ are described below:
[…]
Here $\epsilon_{1}$ and $\epsilon_{2}$ are two positive numbers up to the domain $\Omega$ (see Lemma 2.1).
The linear elliptic PDE in non-divergence form is find the solution $u$ satisfying
(2) $A:D^{2}u+\boldsymbol{b}\cdot\nabla u+cu=f\text{ in }\Omega,\qquad u=0\text{ on }\partial\Omega.$
We assume that $A=A(\boldsymbol{x})\in[L^{\infty}(\Omega)]^{d\times d}$ is a symmetric and uniformly positive definite matrix-valued function on $\Omega$, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, and $c\in L^{\infty}(\Omega)$ with $c\leq 0$.
The elliptic Hamilton-Jacobi-Bellman (HJB) equation is to find the solution $u$ satisfying
(3) $\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}u+\boldsymbol{b}^{\alpha}\cdot\nabla u+c^{\alpha}u-f^{\alpha}]=0\text{ in }\Omega,\qquad u=0\text{ on }\partial\Omega.$
Here the non-empty set $\Lambda$ is called index set. We assume that there are positive constants $\nu\leq\overline{\nu}$ such that for any $\boldsymbol{\xi}\in\mathbb{R}^{d}$ and any $\alpha\in\Lambda$,
(4) $\nu|\boldsymbol{\xi}|^{2}\leq\Sigma_{1\leq i,j\leq d}a_{ij}^{\alpha}(\boldsymbol{x})\xi_{i}\xi_{j}\leq\overline{\nu}|\boldsymbol{\xi}|^{2}\text{ and }c^{\alpha}(\boldsymbol{x})\leq 0,\quad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere};$
(5) $\sup_{\alpha\in\Lambda}\|\boldsymbol{b}^{\alpha}\|_{L^{\infty}(\Omega)}+\sup_{\alpha\in\Lambda}\|c^{\alpha}\|_{L^{\infty}(\Omega)}<+\infty.$
We further assume that there is a subset $E_{0}\subset\Omega$ with zero $d$-dimensional Lebesgue measure such that
(6) $\text{there is a countable subset }\tilde{\Lambda}\subset\Lambda\text{ satisfying that}$
[…]
$\qquad|A^{\alpha}(\boldsymbol{x})-A^{\tilde{\alpha}}(\boldsymbol{x})|+|\boldsymbol{b}^{\alpha}(\boldsymbol{x})-\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|+|c^{\alpha}(\boldsymbol{x})-c^{\tilde{\alpha}}(\boldsymbol{x})|+|f^{\alpha}(\boldsymbol{x})-f^{\tilde{\alpha}}(\boldsymbol{x})|<\epsilon.$
Here, we denote by $|B|=\left(\Sigma_{1\leq i,j\leq d}|b_{ij}|^{2}\right)^{\frac{1}{2}}$ for any matrix $B\in\mathbb{R}^{d\times d}$.
The linear elliptic PDE in non-divergence form (2) is the starting point to study the elliptic HJB equation (3), because (3) becomes (2) if the index set $\Lambda$ contains a single element. In addition, (2)appears in the linearization and numerical methods of fully non-linear second order PDEs [5, 16, 31]. The HJB equation (3) provides a fundamental characterization of the value functions associated with stochastic control problems, which have widespread applications across engineering, physics, economics, and finance [19, 27].
1.1. Literature review
Unlike partial differential equations expressed in divergence form, the theory governing linear equations in non-divergence form, and more broadly, fully nonlinear PDEs, relies on alternative solution frameworks, including classical solutions, viscosity solutions, and strong solutions.
There are several papers [1, 3, 6, 8, 12, 15, 19, 27, 29, 33, 32, 37] focusing on convergence to the viscosity solutions. In the Barles–Souganidis framework [1], monotonicity serves as a fundamental principle underpinning numerical schemes to ensure convergence to viscosity solutions of fully nonlinear partial differential equations. This concept is also applicable to the monotone finite difference methods for the HJB equation [3, 19]. A monotone finite element like scheme for the HJB equation was proposed in [6]. In [32], Nochetto and Zhang introduced a two-scale method based on the integro-differential formulation of (2), where a discrete Alexandroff–Bakelman–Pucci estimate was established. This two-scale method [32] was extended to the Isaacs equation [37]. However, up to our knowledge, all these numerical methods for convergence to the viscosity are low order. In addition, most of them have stencils which increase under mesh refinement, such that the corresponding linear system may not be sparse.
With respect to numerical methods for the convergence to the strong solutions, there are tree main categories.
  • (I)
    AA can be discontinuous but $\gamma A$ dominated by $I_{d}$ where $\gamma$ is a positive weight function. In [39], authors introduced and analyzed a Discontinuous Galerkin (DG) method for (2). The main assumptions are the coefficients $A$ is dominated by $I_{d}$ after being multiplied by a positive weight function $\gamma$ (Cordès condition) and the domain $\Omega$ is convex. The stability and optimal rate of convergence in discrete $H^{2}$-norm were proven in [39]. Following this direction, there are several papers [2, 21, 43]. The technique in [39] has been extended to the HJB equation (3) in [40, 22, 43] but with assumption that $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space. It seems that numerical methods along this direction may only be able to obtain optimal rate of convergence with respect to discrete $H^{2}$-type norm ($p=2$).
  • (II)
    AA is globally continuous. In [14], authors introduced and analyzed a $C^{0}$-conforming DG method for (2). The main assumptions are $A\in C^{0}(\overline{\Omega})$ and the domain $\Omega$ has $C^{1,1}$ boundary. They followed the approach in [24, Chapter $9$] to prove that their numerical method could achieve stability and optimal rate of convergence in discrete $W^{2,p}$-norm for any $1<p<+\infty$. However, this numerical method and its variant - a DG method in [17] can not deal with discontinuous coefficients $A$, since they have the jump of $A\nabla u_{h}$ along mesh interfaces ($u_{h}$ is the numerical solution). Therefore, the numerical methods in [14, 17] can not be applied for the HJB equation (3).
  • (III)
    Residual minimization methods. The main advantage of residual minimization methods [35, 41, 23] is that they can easily inherit the a priori estimate of PDE operators of (2) and (3). However, the minimal requirements of the regularity of true solution $u$ for these methods are highly affected by the choice of the norm to measure the residual in their designs. It is not clear which norm to choose for the two scenarios mentioned above.
A natural question is whether it is possible to design one numerical method for both linear elliptic PDE in non-divergence form and the elliptic HJB equation on Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and achieve optimal convergence in discrete $W^{2,p}$-norm to the strong solution with a wide range of $p$.
1.2. Main contributions
We introduce the finite element methods (14) and (16) for the linear elliptic PDE in non-divergence form (2) and the HJB equation (3), respectively. It is easy to see that the FEM (16) becomes (14), if the index set $\Lambda$ has a single element. Therefore, we can claim that essentially, there is only one FEM for (2) and (3). We would like to point out that the discrete gradient operator $\overline{\nabla}_{h}$ (see (11)) used in (14) and (16) has been introduced in [2, 11, 18].
1.2.1. Contributions for (2)
For the linear elliptic PDE in non-divergence form (2), the main theoretical results are Theorem 3.4, Theorem 3.5 and Theorem 3.6.
Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A\in C^{0}(\overline{\Omega})$. The results of well-posedness of (2) provided by [24, Theorem $9.15$] and optimal convergence of the FEM in [14] require $\partial\Omega$ is $C^{1,1}$. Actually, in neighbourhood of any point on $\partial\Omega$, the analysis in [24, Theorem $9.15$] eventually relies on [24, Lemma $9.12$]. The proof of [24, Lemma $9.12$] needs to “flatten” $\partial\Omega$ such that $\partial\Omega$ needs to be $C^{1,1}$. In contrast, Theorem 3.4 is valid on Lipschitz polyhedral domains (see (1) for the detailed description). Furthermore, the existence of strong solution is mainly proven by the discrete compactness - Lemma 3.8 and the stability of the FEM - Theorem 3.3, while the convergence analysis [14, Theorem 3.2] is based on [24, Theorem $9.15$]. Therefore, even for (2) with uniformly continuous coefficient matrix $A$, the numerical analysis in this paper is quite different from that in [14]. We refer to the statement at the beginning of Section 3 for detailed description of the methodologies used in this paper for (2) with uniformly continuous $A$.
Theorem 3.5 and Theorem 3.6 provides well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$, where the positive weight function $\gamma$ is introduced in (12). Since Theorem 3.5 and Theorem 3.6 are special examples of Theorem 4.1 and Theorem 4.3 respectively, the comparison below between Theorem 3.5, Theorem 4.3 and results in [40] for the HJB equation (3) is directly relevant to Theorem 3.5 and Theorem 3.6 as well.
1.2.2. Contributions for (3)
For the elliptic HJB equation (3), the main theoretical results are Theorem 4.1 and Theorem 4.3. With (4,5) and the assumption (6), Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (3), while Theorem 4.3 provides optimal convergence of the FEM (16). Please see Remark 4.2 and Remark 4.4 for the special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$.
When the domain is convex and $p=2$, the condition (89) used in Theorem 4.3 is more restrictive than the Cordès condition used in [40]. It is due to two reasons. The first reason is that the constant of the discrete Miranda–Talenti inequality of the discrete Lapalcian operator $L_{I,h}$ associated with the FEM (16) is not known explicitly and probably bigger than $1$. The second reason is that we don’t want to include the parameter $\lambda$, which may not be known beforehand, in the design of the weight function $\gamma^{\alpha}$ (see the difference between (15) and the weight function used in [40]). Though it is more restrictive, Remark 4.5 shows that the condition (89) allows non-trivial examples. On the other hand, Theorem 4.3 is valid for a wide range of $p$ (see (1b)) and is valid on two dimensional non-convex polygon for a neighbourhood of $\frac{4}{3}$.
We would like to emphasize that [40] assumes $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space, while [23, Proposition $4.3$] assumes that there are $\beta\in(0,1]$ and $0<C<+\infty$ such that $\|A^{\alpha}\|_{C^{0,\beta}(\overline{\Omega})}<C$ for any $\alpha\in\Lambda$ ([23, Proposition $4.3$] assumes $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$, $c^{\alpha}=0$, and $f^{\alpha}=f$). So, if the index set $\Lambda$ has a single element, the elliptic HJB equation (3) will not become the linear elliptic PDE in non-divergence form (2) with discontinuous matrix coefficient $A$. In contrast, the assumption (6) allows discontinuity of $A^{\alpha}$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. By Lemma 4.3, we can conclude that the assumptions used by [40] and ([23, Proposition $4.3$] imply that the assumption (6) holds.
If the index set $\Lambda$ is countable, then for any $w\in W^{2,1}(\Omega)$, the function
$\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]$
is a measurable function on $\Omega$. However, it is not obviously true if $\Lambda$ is uncountable. Lemma 4.1 shows that the assumption (6) implies (see (74))
$\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=\sup_{\tilde{\alpha}\in\tilde{\Lambda}}[A^{\tilde{\alpha}}:D^{2}w+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}$],
where $\tilde{\Lambda}$ is a countable subset of $\Lambda$. Thus this issue is solved with the assumption (6) which is more general than assumptions in [40] and [23, Proposition $4.3$].
1.3. Structure of this paper
The rest of this paper is organized as follows. In Section 22, we provide basic notations, the FEMs for (2) and (3), and some preliminary results. In Section 33, we provide theoretical results for (2). In Section 44, we provide theoretical results for (3).
2. Assumptions, notations, FEM and preliminary results
In Section 2.1, we provide the assumption of mesh, the discrete gradient operator (11) and several basic notations. In Section 2.2, we introduce the FEM (14) and the FEM (16) for (2) and (3), respectively. Actually, the FEM (16) becomes the FEM (14) if the index set $\Lambda$ has a single element. In Section 2.3, we provide some preliminary results.
2.1. Mesh and space notation
We denote by $\mathcal{T}_{h}$ a quasi-uniform, simplicial, and conforming triangulation of the domain $\Omega$. Let $\mathcal{F}_{h}^{I}$ be the collection of all (d1)(d-1)-dimensional interior mesh faces in $\mathcal{T}_{h}$, $\mathcal{F}_{h}^{B}$ the collection of (d1)(d-1)-dimensional boundary mesh faces in $\mathcal{T}_{h}$, and $\mathcal{F}_{h}:=\mathcal{F}_{h}^{I}\cup\mathcal{F}_{h}^{B}$ the collection of all (d1)(d-1)-dimensional mesh faces in $\mathcal{T}_{h}$.
Let $K^{+},K^{-}\in\mathcal{T}_{h}$ and $F=\partial K^{+}\cap\partial K^{-}\in\mathcal{F}_{h}^{I}$. Without loss of generality, we assume that the global labelling number of $K^{+}$ is larger than that of $K^{-}$. We introduce the jump and average of a scalar or vector valued function $v$ as
$[v]|_{F}:=v^{+}-v^{-},\qquad\{v\}|_{F}=\frac{1}{2}\left(v^{+}+v^{-}\right),$
where $v^{\pm}=v|_{K^{\pm}}$. On a boundary mesh face $F\in\mathcal{F}_{h}^{B}$ with $F=\partial K^{+}\cap\partial\Omega$, we define $v]|_{F}=\{v\}|_{F}=v^{+}$.
We define $V_{h}=H_{0}^{1}(\Omega)\cap P_{r}(\mathcal{T}_{h})$ and $\overline{V}_{h}=P_{r}(\mathcal{T}_{h})$ with integer $r\geq 2$. We also define the piecewise Sobolev space with respect to the mesh $\mathcal{T}_{h}$:
$W^{s,p}(\mathcal{T}_{h}):=\boldsymbol{\Pi}_{K\in\mathcal{T}_{h}}W^{s,p}(K),\qquad W_{h}^{p}:=W^{2,p}(\mathcal{T}_{h})\cap W_{0}^{1,p}(\Omega),\qquad W_{h}^{s,p}(D):=W^{s,p}(\mathcal{T}_{h})|_{D}.$
For a given subdomain $D\subset\Omega$. we define $V_{h}(D)\subset V_{h}$ and $W_{h}^{p}(D)\subset W_{h}^{p}(D)$ as the subspaces which vanish outside of $D$ by
$V_{h}(D):=\{v\in V_{h}:v_{h}|_{\Omega\setminus D}=0\},\qquad W_{h}^{p}(D):=\{v\in W_{h}^{p}:v|_{\Omega\setminus D}=0\}.$
Associated with $D\subset\Omega$, we define a semi-norm on $W_{h}^{2,p}(D)$ for $1<p<\infty$:
(7) $\|v\|_{W_{h}^{2,p}(D)}=\|D_{h}^{2}v\|_{L^{p}(D)}+\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\nabla v]\|_{L^{p}(F\cap\overline{D})}^{p}\right)^{\frac{1}{p}}.$
Here $D_{h}^{2}v$ denotes the element-wise Hessian matrix of $v$. Obviously,
$\|v\|_{W_{h}^{2,p}(\Omega)}=\|D_{h}^{2}v\|_{L^{p}(\Omega)}+\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\nabla v]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}.$
For any subdomain $D\subset\Omega$ whose inner radius bigger than $2h$ and any $w\in L_{h}^{p}(D)$, we introduce the following mesh-dependent semi-norm
(8) $\|w\|_{L_{h}^{p}(\Omega)}:=\sup_{v_{h}\in V_{h}(D)}\dfrac{(w,v_{h})_{D}}{\|v_{h}\|_{L^{p^{\prime}}(D)}},$
where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
For any $F=\partial K^{+}\cap\partial K^{-}\in\mathcal{F}_{h}^{I}$, we denote by $\boldsymbol{n}_{F}$ the unit outward normal vector pointing in the direction of the element with the smaller global index. Therefore $\boldsymbol{n}_{F}=\boldsymbol{n}_{K^{+}}=-\boldsymbol{n}_{K^{-}}$. For $F\in\mathcal{F}_{h}^{B}$, we define $\boldsymbol{n}_{F}$ to be the unit outward normal vector to $\partial\Omega$ restricted to $F$.
Definition 2.1.
We denote by $W^{1,1}(\mathcal{T}_{h}):=\{\phi\in L^{1}(\Omega):\phi|_{K}\in W^{1,1}(K),\forall K\in\mathcal{T}_{h}\}$. We define the discrete partial derivatives $\overline{\partial}_{h,x_{i}}:W^{1,1}(\mathcal{T}_{h})\rightarrow\overline{V}_{h}$ by
(9) $(\overline{\partial}_{h,x_{i}}v,\varphi_{h})_{\mathcal{T}_{h}}=\left\langle\{v\}n^{(i)},[\varphi_{h}]\right\rangle_{\mathcal{F}_{h}}-(v,\partial_{x_{i}}\varphi_{h})_{\mathcal{T}_{h}},\qquad\forall\varphi_{h}\in\overline{V}_{h}.$
Here $v\in W^{1,1}(\mathcal{T}_{h})$, $i\in\{1,\cdots,d\}$ and $\boldsymbol{n}_{F}=(n_{F}^{(1)},\cdots,n_{F}^{(d)})^{\top}$ for any $F\in\mathcal{F}_{h}$. It is easy to see that
(10) $(\overline{\partial}_{h,x_{i}}v,\varphi_{h})_{\mathcal{T}_{h}}=-\left\langle[v]n^{(i)},\{\varphi_{h}\}\right\rangle_{\mathcal{F}_{h}^{I}}+(\partial_{x_{i}}v,\varphi_{h})_{\mathcal{T}_{h}},\qquad\forall\varphi_{h}\in\overline{V}_{h}.$
The discrete gradient operator $\overline{\nabla}_{h}:W^{1,1}(\mathcal{T}_{h})\rightarrow[\overline{V}_{h}]^{d}$ is defined as
(11) $\left(\overline{\nabla}_{h}v\right)_{i}=\overline{\partial}_{h,x_{i}}v,\qquad\forall i\in\{1,\cdots,d\}.$
We would like to point out that the discrete gradient operator (11) has been introduced in [2, 11, 18].
2.2. Finite element method for linear elliptic PDE in non-divergence form and HJB equation
2.2.1. FEM for linear elliptic PDE in non-divergence form
We define a function $\gamma$ on $\Omega$ to be
(12) $\gamma(\boldsymbol{x})=\dfrac{\text{Tr}A(\boldsymbol{x})}{|A(\boldsymbol{x})|^{2}}=\dfrac{\Sigma_{1\leq i\leq d}a_{ii}(\boldsymbol{x})}{\Sigma_{1\leq i,j\leq d}(a_{ij}(\boldsymbol{x}))^{2}},\qquad\forall\boldsymbol{x}\in\Omega.$
Obviously, the linear elliptic PDE in non-divergence form (2) is equivalent to
(13) $\tilde{A}:D^{2}u+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u=\tilde{f}\text{ in }\Omega,\qquad u=0\text{ on }\partial\Omega,$
where $\tilde{A}=\gamma A$, $\tilde{\boldsymbol{b}}=\gamma\boldsymbol{b}$, $\tilde{c}=\gamma c$ and $\tilde{f}=\gamma f$.
The finite element method for linear elliptic PDE in non-divergence form (2) is to find $u_{h}\in V_{h}$ satisfying
(14) $\int_{\Omega}(\tilde{A}:\overline{\nabla}_{h}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u_{h}+\tilde{c}u_{h})v_{h}d\boldsymbol{x}$
$=$ $\int_{\Omega}\gamma(\tilde{A}:\overline{\nabla}_{h}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u_{h}+\tilde{c}u_{h})v_{h}d\boldsymbol{x}=\int_{\Omega}\gamma fv_{h}d\boldsymbol{x}=\int_{\Omega}\tilde{f}v_{h}d\boldsymbol{x},\qquad\forall v_{h}\in V_{h}.$
2.2.2. FEM for the Hamilton-Jacobi-Bellman equation
For any $\alpha\in\Lambda$, we define a function $\gamma^{\alpha}$ on $\Omega$ to be
(15) $\gamma^{\alpha}(\boldsymbol{x})=\dfrac{\text{Tr}A^{\alpha}(\boldsymbol{x})}{|A^{\alpha}(\boldsymbol{x})|}=\dfrac{\Sigma_{1\leq i\leq d}a_{ii}^{\alpha}(\boldsymbol{x})}{\Sigma_{1\leq i,j\leq d}(a_{ij}^{\alpha}(\boldsymbol{x}))^{2}},\qquad\forall\boldsymbol{x}\in\Omega.$
The finite element method for the Hamilton-Jacobi-Bellman equation (3) is to find $u_{h}\in V_{h}$ satisfying
(16) $(\sup_{\alpha\in\Lambda}\gamma^{\alpha}[A^{\alpha}:\overline{\nabla}_{h}(\nabla u_{h})+\boldsymbol{b}^{\alpha}\cdot\nabla u_{h}+c^{\alpha}u_{h}-f^{\alpha}],v_{h})_{\Omega}=0,\qquad\forall v_{h}\in V_{h}.$
2.3. Preliminary results
Lemma 2.1.
If the domain $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), let $p\in(1,2]$. If the domain is a Lipschitz polygon in $\mathbb{R}^{2}$, let $p\in(1,\frac{4}{3}+\epsilon_{2})$ where $\epsilon_{2}>0$ depends only on $\Omega$. Then for any $g\in L^{p}(\Omega)$, there is a unique solution $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ such that $\Delta w=g$ in $\Omega$ and
(17) $\|w\|_{W^{2,p}(\Omega)}\leq\underline{C}_{p}\|g\|_{L^{p}(\Omega)}.$
Here the constant $\underline{C}_{p}$ may depend on $p$.
If the domain $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), let $p\in(1,+\infty)$. If the domain is a Lipschitz polygon in $\mathbb{R}^{2}$, let $p\in(\frac{4}{3}-\epsilon_{1},3+\epsilon_{3})$ where $\epsilon_{1},\epsilon_{3}>0$ depend only on $\Omega$. Then for any $g\in W^{-1,p}(\Omega)$, there is a unique solution $w\in W_{0}^{1,p}(\Omega)$ such that $\Delta w=g$ in $\Omega$ and
(18) $\|w\|_{W^{1,p}(\Omega)}\leq\underline{C}_{p}^{\prime}\|g\|_{W^{-1,p}(\Omega)}.$
Here the constant $\underline{C}_{p}^{\prime}$ may depend on $p$.
In addition, if $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and $g\in L^{2}(\Omega)$, we have
(19) $\|D^{2}w\|_{L^{2}(\Omega)}\leq\|g\|_{L^{2}(\Omega)}.$
Proof.
The proof of existence of a unique solution and the estimate (17) are due to [7, Proposition $2.1$] ($d=2$) and [20, ($1.6$)] ($d=3$). The proof of existence of a unique solution and the estimate (18) are due to [28, Theorem $0.5$] (non-convex and $d=2$) and [20, ($1.5$)] (convex). The estimate (19) is due to [39, Lemma $1$]. ∎
According to [9, Theorem $3$ and Theorem $4$] (see [42, Lemma $7.3$] as well), we have the following Lemma 2.2 since we assume the mesh $\mathcal{T}_{h}$ is quasi-uniform. We would like to point out that the proof of [9, Theorem $4$] can be easily extended to obtain (20b) in three dimensional domains.
Lemma 2.2.
We denote by $P_{h}$ the standard $L^{2}$-orthogonal projection onto $V_{h}$. Then there is a constant $\tilde{C}_{1}$ such that for any $q\in[1,+\infty)$,
(20a) $\|P_{h}v\|_{L^{q}(\Omega)}\leq\tilde{C}_{1}\|v\|_{L^{q}(\Omega)}\qquad\forall v\in L^{q}(\Omega),$
(20b) $\|P_{h}v\|_{W^{1,q}(\Omega)}\leq\tilde{C}_{1}\|v\|_{W^{1,q}(\Omega)}\qquad\forall v\in W_{0}^{1,p}(\Omega).$
According to [13, Lemma $4.3$], we have the following Lemma 2.3.
Lemma 2.3.
For any $\overline{v}_{h}\in\overline{V}_{h}$, we denote by $\chi_{h}\in H^{1}(\Omega)\cap\overline{V}_{h}$ the standard $H^{1}$-conforming interpolation of $\overline{v}_{h}$. Then for any $q\in[1,+\infty]$, we have
$\|\chi_{h}-\overline{v}_{h}\|_{L^{q}(\Omega)}\leq Ch\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-q}\|[\overline{v}_{h}]\|_{L^{q}(F)}^{q}\right)^{\frac{1}{q}}.$
3. Linear elliptic PDE in non-divergence form
For the linear elliptic PDE in non-divergence form (2), the main theoretical results are Theorem 3.4, Theorem 3.5 and Theorem 3.6.
Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A\in C^{0}(\overline{\Omega})$. We explain the methodology to prove Theorem 3.4 by the following steps.
  • Step $1$. We prove Theorem 3.2 which give the following global $W^{2,p}$ estimate (22) of the linear elliptic PDE in non-divergence form (2). The most difficult part of the proof of Theorem 3.2 is to obtain (22) for $1<p\leq\frac{3}{2}$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$, since the Alexandroff–Bakelman–Pucci estimate can not be applied immediately due to the fact $W^{2,p}(\Omega)$ can not be embedded into $C^{0}(\overline{\Omega})$. To overcome this difficulty, we provide Lemma 3.1 which shows the uniqueness of the solution of (2) in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$.
  • Step $2$. We provide Lemma 3.7 which give the inequality (38). The proof of Lemma 3.7 mimics that of [14, Lemma $3.4$]. However, due to the construction of the discrete gradient operator (11), the proof of Lemma 3.7 is much more technical.
  • Step $3$. To erase the term $\|w_{h}\|_{W^{1,p}(\Omega)}$ on the right hand side of (38), we prove by contradiction with the help of Lemma 3.8 which provides the discrete compactness. Then we obtain Theorem 3.2 which the stability of the FEM (14).
  • Step $4$. Roughly speaking, we apply the discrete compactness provided by Lemma 3.8 with Theorem 3.2 to prove the limit of numerical solution of the FEM (14) is a solution of (2). Then by the global $W^{2,p}$ estimate (22), the well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) is proven for the range of $p$ described in (1b). Finally by some density argument, we extend $p$ to the range described in (1a). Optimal convergence of the numerical solution of the FEM (14) is straightforward.
Theorem 3.5 and Theorem 3.6 provides well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$, where the positive weight function $\gamma$ is introduced in (12). Since they are special examples of Theorem 4.1 and Theorem 4.3 respectively, the proofs of Theorem 3.5 and Theorem 3.6 are skipped. Please refer to those of Theorem 4.1 and Theorem 4.3.
3.1. Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form with uniformly continuous AA
Lemma 3.1.
(Uniqueness of linear elliptic PDE in non-divergence form) Let $A\in[C^{0}(\overline{\Omega})]^{3\times 3}$ uniformly elliptic, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{3}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ is an open bounded convex domain in $\mathbb{R}^{3}$. Let $1<p\leq\frac{3}{2}$. If $\phi\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ such that
$A:D^{2}\phi+\boldsymbol{b}\cdot\nabla\phi+c\phi=0\text{ in }\Omega,$
then $\phi=0$ in $\Omega$.
Remark 3.1.
Though it is inspired by that of [24, Lemma $9.16$], the proof of Lemma 3.1 avoids using the common technique of “flattening” $\partial\Omega$, which is critical in [24, Lemma $9.16$]. Thus Lemma 3.1 is valid on convex polyhedra in $\mathbb{R}^{3}$.
Proof.
We choose $\boldsymbol{x}_{0}\in\overline{\Omega}$ arbitrarily. For any $R>0$, we denote by $\eta$ the cut-off function to be $1$ inside of $\mathcal{B}_{R}(\boldsymbol{x}_{0}):=\{\boldsymbol{x}\in\mathbb{R}^{3}:|\boldsymbol{x}-\boldsymbol{x}_{0}|<R\}$ and to be 0 outside of $\mathcal{B}_{2R}(\boldsymbol{x}_{0}):=\{\boldsymbol{x}\in\mathbb{R}^{3}:|\boldsymbol{x}-\boldsymbol{x}_{0}|<2R\}$.
We define $v=\eta\phi$. We notice that in $\Omega$,
$A:D^{2}v=A:D^{2}(\eta\phi)$
$=$ $\eta A:D^{2}\phi+2A:(\nabla\eta\otimes\nabla\phi)+\phi A:D^{2}\eta$
$=$ $-\eta\left(\boldsymbol{b}\cdot\nabla\phi+c\phi\right)+2A:(\nabla\eta\otimes\nabla\phi)+\phi A:D^{2}\eta,$
since $A:D^{2}\phi+\boldsymbol{b}\cdot\nabla\phi+c\phi=0$ in $\Omega$. Therefore, by Sobolev’s embedding inequality, we have that
(21) $A:D^{2}v\in L^{q}(\Omega)\text{ with }q:=\min(\dfrac{3p}{3-p},2)>\frac{3}{2}.$
Obviously, $1<p<q$.
We define $A_{0}=A(\boldsymbol{x}_{0})\in\mathbb{R}^{3\times 3}$. We define the linear mapping $M:L^{\tilde{q}}(\Omega)\rightarrow W^{2,\tilde{q}}(\Omega)\cap W_{0}^{1,\tilde{q}}(\Omega)$ by
$-\nabla\cdot\left(A_{0}\nabla M(f)\right)=f,\qquad\forall f\in L^{\tilde{q}}(\Omega),\forall 1<\tilde{q}\leq q.$
By Lemma 2.1, the mapping $M$ is not only well-defined but also continuous for any $1<\tilde{q}\leq q$.
It is easy to see that in $\Omega$
$A_{0}:D^{2}v=(A_{0}-A):D^{2}v+A:D^{2}v.$
Since $1<p<q$, we can apply the mapping $M$ on both sides of the above equation, to obtain that
$v=M\left((A_{0}-A):D^{2}v\right)+M\left(A:D^{2}v\right)\text{ in }\Omega.$
We define $B_{2R}(\boldsymbol{x}_{0}):=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{0}|<2R\}=\mathcal{B}_{2R}(\boldsymbol{x}_{0})\cap\Omega$. We define a linear mapping $T:W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))\rightarrow W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))$ by
$Tw=\left(M\tilde{w}\right)|_{B_{2R}(\boldsymbol{x}_{0})}$
where $w\in W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))$ and
$\tilde{w}=(A_{0}-A):D^{2}w\text{ in }B_{2R}(\boldsymbol{x}_{0}),\qquad\tilde{w}=0\text{ in }\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}.$
Here $1<\tilde{q}\leq q$. Therefore, for any $1<\tilde{q}\leq q$, we have that
$\|Tw\|_{W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))}\leq C_{\tilde{q}}\|\tilde{w}\|_{L^{\tilde{q}}(\Omega)}$
$=$ $C_{\tilde{q}}\|(A_{0}-A):D^{2}w\|_{L^{\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))}$
$\leq$ $C_{\tilde{q}}\|A_{0}-A\|_{L^{\infty}(B_{2R}(\boldsymbol{x}_{0}))}\|w\|_{W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0}))},\qquad\forall w\in W^{2,\tilde{q}}(B_{2R}(\boldsymbol{x}_{0})).$
The constant $C_{\tilde{q}}$ is due to Lemma 2.1.
Since $A\in[C^{0}(\overline{\Omega})]^{3\times 3}$, there is a $R_{0}>0$ such that $\|T\|\leq\frac{1}{2}$ from $W^{2,\tilde{q}}(B_{2R_{0}}(\boldsymbol{x}_{0}))$ to $W^{2,\tilde{q}}(B_{2R_{0}}(\boldsymbol{x}_{0}))$ for all $p\leq\tilde{q}\leq q$. By (21) and Lemma 2.1, $M\left(A:D^{2}v\right)\in L^{\tilde{q}}(\Omega)$ for any $1<\tilde{q}\leq q$. Therefore for any $p\leq\tilde{q}\leq q$, there is a unique $w\in W^{2,\tilde{q}}(B_{2R_{0}}(\boldsymbol{x}_{0}))$ such that
$w=Tw+M\left(A:D^{2}v\right)|_{B_{2R_{0}(\boldsymbol{x}_{0})}}\text{ in }B_{2R_{0}(\boldsymbol{x}_{0})}.$
On the other hand, since $v=0$ in $\Omega\setminus\overline{B_{2R}(\boldsymbol{x}_{0})}$ and $v\in W^{2,p}(\Omega)$ ($1<p<q$), we have that
$v|_{B_{2R_{0}}(\boldsymbol{x}_{0})}=T(v|_{B_{2R_{0}}(\boldsymbol{x}_{0})})+M\left(A:D^{2}v\right)|_{B_{2R_{0}}(\boldsymbol{x}_{0})}\text{ in }B_{2R_{0}}(\boldsymbol{x}_{0}).$
Therefore we have that $\phi|_{B_{R_{0}}(\boldsymbol{x}_{0})}\in W^{2,q}(B_{R_{0}}(\boldsymbol{x}_{0}))$.
Since $\boldsymbol{x}_{0}\in\overline{\Omega}$ is chosen arbitrarily, we can conclude that $\phi\in W^{2,q}(\Omega)$ where $q=\min(\dfrac{3p}{3-p},2)>\frac{3}{2}$. So $\phi\in C^{0}(\overline{\Omega})$. Furthermore, by the interior estimate in [24, Lemma $9.16$], we have $\phi\in W_{\text{loc}}^{2,3}(\Omega)$ since $A:D^{2}\phi+\boldsymbol{b}\cdot\nabla\phi+c\phi=0$ in $\Omega$. Therefore, by the ABP estimate [24, Theorem $9.1$], we can conclude that $\phi=0$ in $\Omega$. ∎
Theorem 3.2.
(Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form) Let $A\in[C^{0}(\overline{\Omega})]^{d\times d}$, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ is a bounded open Lipschitz polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). We further assume $A$ is uniformly elliptic in $\Omega$. There is a constant $C_{p}>0$ such that
(22) $\|w\|_{W^{2,p}(\Omega)}\leq C_{p}\|A:D^{2}w+\boldsymbol{b}\cdot\nabla w+cw\|_{L^{p}(\Omega)},\qquad\forall w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega).$
Here $p$ is valid in the range described in (1a).
Proof.
When $d=2$, $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ due to Sobolev’s embedding theorem. Furthermore, by the interior estimate in [24, Lemma $9.16$], we have $u\in W_{\text{loc}}^{2,2}(\Omega)$. Therefore, Theorem 3.2 can be proven by mimicking the proofs of [34, Theorem $1.1$] and [34, Corollary $1.2$] with the global $W^{2,p}$ estimate of Poisson equation (see [7, Proposition $2.1$]) and the Alexandroff–Bakelman–Pucci estimate estimate [24, Theorem $9.1$].
Now we focus on the case $d=3$. Due to [34, Corollary $1.2$], Theorem 3.2 holds if we further assume $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$ and $\frac{3}{2}<p\leq 2$. In fact, if we replace [10, Corollary $3.12$] by Lemma 2.1 in the proof of [34, Theorem $1.1$], it is easy to have that for any $1<p\leq 2$ and for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
(23) $\|w\|_{W^{2,p}(\Omega)}\leq C_{p}^{\prime}\left(\|A:D^{2}w+\boldsymbol{b}\cdot\nabla w+cw\|_{L^{p}(\Omega)}+\|w\|_{L^{p}(\Omega)}\right).$
Here $C_{p}^{\prime}$ depends on $1<p\leq 2$ and $\Omega$ is an open bounded convex domain in $\mathbb{R}^{3}$.
As the proof of [34, Corollary $1.2$], Theorem 3.2 will be proven via (23) and the proof of contraction with the uniqueness of solution in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$:
$A:D^{2}w+\boldsymbol{b}\cdot\nabla w+cw=0\text{ in }\Omega\text{ with }w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)\Longrightarrow u=0\text{ in }\Omega.$
When $\frac{3}{2}<p\leq 2$, it is straightforward by the Alexandroff–Bakelman–Pucci estimate estimate with the fact $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ (see the proof of [34, Corollary $1.2$]). Therefore, in order to complete the proof of Theorem 3.2, we only need the uniqueness for $1<p\leq\frac{3}{2}$, which is provided by Lemma 3.1. Therefore, we can conclude that the proof of Theorem 3.2 is complete. ∎
3.2. Stability analysis of FEM for linear elliptic PDE in non-divergence form with uniformly continuous AA
3.2.1. Stability estimates for the linear elliptic PDE in non-divergence form with constant $A$ and $\boldsymbol{b}=\boldsymbol{0},c=0$
All results in Section 3.2.1 can be found in [14]. But the results in [14] are only valid on domains with $C^{1,1}$ boundary.
We denote by $\tilde{A}_{0}\in\mathbb{R}^{d\times d}$ a symmetric and positive definite matrix.
According to Definition 2.1 and the fact $V_{h}\subset\overline{V}_{h}$, we have
(24) $\int_{\Omega}(\tilde{A}_{0}:\overline{\nabla}_{h}(\nabla w_{h}))v_{h}d\boldsymbol{x}=-(\tilde{A}_{0}\nabla w_{h},\nabla v_{h})_{\Omega},\qquad\forall w_{h},v_{h}\in V_{h}.$
We define a linear operator $\mathcal{L}_{\tilde{A}_{0},h}:\rightarrow W_{h}^{1}\rightarrow V_{h}$ as
(25) $(\mathcal{L}_{\tilde{A}_{0},h}w,v_{h})_{\Omega}=\int_{\Omega}(\tilde{A}_{0}:\overline{\nabla}_{h}(\nabla w))v_{h}d\boldsymbol{x},\qquad\forall w\in W_{h}^{1},v_{h}\in V_{h}.$
Here, we recall that the piecewise Sobolev space $W_{h}^{1}=W^{2,1}(\mathcal{T}_{h})\cap W_{0}^{1,1}(\Omega)$.
Lemma 3.2.
There is a positive constant $C$ such that for any $h\in(0,h_{0})$,
(26) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\sup_{\chi_{h}\in V_{h}}\|\chi_{h}-w\|_{W^{1,p}(\Omega)},\qquad\forall w\in W_{0}^{1,p}(\Omega).$
Here $w_{h}\in V_{h}$ satisfying
$(\tilde{A}_{0}\nabla w_{h},v_{h})_{\Omega}=(\tilde{A}_{0}\nabla w,v_{h})_{\Omega},\qquad\forall v_{h}\in V_{h}.$
If the domain $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), the estimate (26) holds for any $p\in(1,+\infty]$. If the domain is a Lipschitz polygon in $\mathbb{R}^{2}$, the estimate (26) holds for any $p\in(\frac{4}{3}-\epsilon_{1},+\infty]$ where $\epsilon_{1}$ is the same as that in Lemma 2.1. The constants $C$ and $h_{0}$ may depend $\tilde{A}_{0}$.
Proof.
By [25, Theorem $2$] ($d=3$), [38, Theorem $2$] ($d=2$),
(27) $\|\nabla(w_{h}-w)\|_{L^{\infty}(\Omega)}\leq C\|\nabla(\chi_{h}-w)\|_{L^{\infty}(\Omega)},\qquad\forall\chi_{h}\in V_{h},$
if $h<h_{0}$. Here the constant $C$ may depend on $\tilde{A}_{0}$. Therefore, there is a constant $C$ such that for any $2\leq p\leq+\infty$, if $h<h_{0}$,
(28) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\|\chi_{h}-w\|_{W^{1,p}(\Omega)},\qquad\forall\chi_{h}\in V_{h}.$
Now, for any $p\in(1,2)$, we denote by $p^{\prime}\geq 2$ such that $\frac{1}{p}+\frac{1}{p^{\prime}}$. We choose $\chi_{h}\in V_{h}$ arbitrarily. Then by Lemma 2.1, we have for any $p\in(1,2)$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and for any $p\in(\frac{4}{3}-\epsilon_{1},2]$ if $\Omega$ is a Lipschitz polygon in $\mathbb{R}^{2}$,
$\|w_{h}-\chi_{h}\|_{W^{1,p}(\Omega)}\leq C\|\nabla\cdot(\tilde{A}_{0}\nabla(w_{h}-\chi_{h}))\|_{W^{-1,p}(\Omega)}$
$=$ $C\sup_{\tilde{v}\in W_{0}^{1,p^{\prime}}(\Omega)}\dfrac{-(\nabla\cdot(\tilde{A}_{0}\nabla(w_{h}-\chi_{h}),\tilde{v}))_{\Omega}}{\|\tilde{v}\|_{W^{1,p^{\prime}}(\Omega)}}=C\sup_{\tilde{v}\in W_{0}^{1,p^{\prime}}(\Omega)}\dfrac{(\tilde{A}_{0}\nabla(w_{h}-\chi_{h}),\nabla\tilde{v})_{\Omega}}{\|\tilde{v}\|_{W^{1,p^{\prime}}(\Omega)}}$
$=$ $C\sup_{\tilde{v}\in W_{0}^{1,p^{\prime}}(\Omega)}\dfrac{(\tilde{A}_{0}\nabla(w_{h}-\chi_{h}),\nabla\tilde{v}_{h})_{\Omega}}{\|\tilde{v}\|_{W^{1,p^{\prime}}(\Omega)}}=C\sup_{\tilde{v}\in W_{0}^{1,p^{\prime}}(\Omega)}\dfrac{(\tilde{A}_{0}\nabla(w-\chi_{h}),\nabla\tilde{v}_{h})_{\Omega}}{\|\tilde{v}\|_{W^{1,p^{\prime}}(\Omega)}},$
where $\tilde{v}_{h}\in V_{h}$ satisfying
$(\tilde{A}_{0}\nabla\tilde{v}_{h},\nabla\tilde{w}_{h})_{\Omega}=(\tilde{A}_{0}\nabla\tilde{v},\nabla\tilde{w}_{h}),\qquad\forall\tilde{w}_{h}\in V_{h}.$
By (28),
$\|\tilde{v}_{h}\|_{W^{1,p^{\prime}}(\Omega)}\leq C\|\tilde{v}\|_{W^{1,p^{\prime}}(\Omega)}.$
So we have
$\|w_{h}-\chi_{h}\|_{W^{1,p}(\Omega)}\leq C\|w-\chi_{h}\|_{W^{1,p}(\Omega)}.$
Therefore, for any $p\in(1,2)$, we also have that if $h<h_{0}$,
(29) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\|\chi_{h}-w\|_{W^{1,p}(\Omega)},\qquad\forall\chi_{h}\in V_{h}.$
(28, 29) imply that for any $p\in(1,+\infty]$,
$\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\inf_{\chi_{h}\in V_{h}}\|\chi_{h}-w\|_{W^{1,p}(\Omega)}.$
Therefore, the proof of the estimates (26) is complete. ∎
Lemma 3.3.
Let $\mathcal{L}_{\tilde{A}_{0},h}$ be the linear operator defined in (25). There is a positive constant $C$ such that for all $h\in(0,h_{0})$,
(30) $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|\mathcal{L}_{\tilde{A}_{0},h}w_{h}\|_{L^{p}(\Omega)},\qquad\forall w_{h}\in V_{h}.$
Here $p$ is valid in the range described in (1b). The constants $C$ and $h_{0}$ may depend $\tilde{A}_{0}$.
Proof.
We choose $w_{h}\in V_{h}$ arbitrarily. We notice that $\mathcal{L}_{\tilde{A}_{0},h}w_{h}\in V_{h}$. We denote by $w\in H_{0}^{1}(\Omega)$ satisfying
$\tilde{A}_{0}:D^{2}w=\nabla\cdot\left(\tilde{A}_{0}\nabla w\right)=\mathcal{L}_{\tilde{A}_{0},h}w_{h}\quad\text{ in }\Omega.$
According to (24),
(31) $(\tilde{A}_{0}\nabla w_{h},\nabla v_{h})_{\Omega}=(\tilde{A}_{0}\nabla w,\nabla v_{h})_{\Omega},\qquad\forall v_{h}\in V_{h}.$
Then by (31), the construction of $w$ and Lemma 3.2, we have that for any $p\in(1,+\infty]$ and $0<h<h_{0}$,
(32) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\inf_{\chi_{h}\in V_{h}}\|\chi_{h}-w\|_{W^{1,p}(\Omega)}\leq Ch\|w\|_{W^{2,p}(\Omega)}.$
(32) and Lemma 2.1 imply that for any $p\in(1,2]$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and for any $p\in(\frac{4}{3}-\epsilon_{1},\frac{4}{3})+\epsilon_{2}$, if $h<h_{0}$,
(33) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq Ch\|w\|_{W^{2,p}(\Omega)}\leq CC^{\prime}h\|\mathcal{L}_{\tilde{A}_{0},h}w_{h}\|_{L^{p}(\Omega)}.$
Here the constant $C^{\prime}$ may also depend on $\tilde{A}_{0}$.
We denote by $\boldsymbol{\phi}_{h}$ the standard $L^{2}$-orthogonal projection of $\nabla w$ into $P_{0}(\mathcal{T}_{h}]^{d}$. Then by the discrete trace inequality, we have
(34) $\|\boldsymbol{\phi}_{h}-\nabla w\|_{L^{p}(\Omega)}\leq Ch\|w\|_{W^{2,p}(\Omega)}\leq CC^{\prime}h\|\mathcal{L}_{\tilde{A}_{0},h}w_{h}\|_{L^{p}(\Omega)},$
$\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\boldsymbol{\phi}_{h}]\|_{L^{p}(F\cap\overline{D})}^{p}\right)^{\frac{1}{p}}\leq C\|w\|_{W^{2,p}(\Omega)}\leq CC^{\prime}\|\mathcal{L}_{\tilde{A}_{0},h}w_{h}\|_{L^{p}(\Omega)}.$
Then by (33,34) and the discrete inverse inequality, we can conclude that the proof is complete. ∎
Lemma 3.4.
For $\boldsymbol{x}_{0}\in\overline{\Omega}$ and $R>0$, define
$B_{R}(\boldsymbol{x}_{0}):=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{0}|<R\}\subset\Omega.$
Let $R^{\prime}=R+d$ with $d\geq 2h$. Then there is a positive constant $C$ such that
(35) $\|w_{h}\|_{W_{h}^{2,p}(B_{R}(\boldsymbol{x}_{0}))}\leq C\|\mathcal{L}_{\tilde{A}_{0},h}w_{h}\|_{L_{h}^{p}(B_{R^{\prime}}(\boldsymbol{x}_{0}))},\qquad\forall w_{h}\in V_{h}(B_{R}(\boldsymbol{x}_{0})),$
for any $h\in(0,h_{0})$ where $h_{0}$ is introduced in Lemma 3.3.
Here $p$ is valid in the range described in (1b). The constant $C$ is independent of the choice of $R$ and $R^{\prime}$. But the constant $C$ may depend on $\tilde{A}_{0}$.
Proof.
By (24), $\mathcal{L}_{\tilde{A}_{0},h}w_{h}$ is the same as $\mathcal{L}_{0,h}$ in [14, Lemma $2.7$]. Therefore, by replacing [14, Lemma $2.6$] by Lemma 3.3, the proof in [14, Lemma $2.7$] is also valid for this Lemma. It is easy to verify that the proof of [14, Lemma $2.7$] is independent of the choice of $\boldsymbol{x}_{0}$, $R$ and $R^{\prime}$. ∎
3.2.2. Stability estimates for linear elliptic PDE in non-divergence form with uniformly continuous AA
Analogues of Lemma 3.5, Lemma 3.6 and Lemma 3.7 can be found in [14].
We define $\mathcal{L}_{\tilde{A},h}:W_{h}^{1}\rightarrow V_{h}$ as
(36) $(\mathcal{L}_{\tilde{A},h}w,v_{h})_{\Omega}=\int_{\Omega}(\tilde{A}:\overline{\nabla}_{h}(\nabla u_{h}))v_{h}d\boldsymbol{x},\qquad\forall v_{h}\in V_{h}.$
Here we recall that $W_{h}^{1}=W_{0}^{1,1}(\Omega)\cap W^{2,1}(\mathcal{T}_{h})$.
For $\boldsymbol{x}_{0}\in\overline{\Omega}$ and $R>0$, we define
$B_{R}(\boldsymbol{x}_{0}):=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{0}|<R\}\subset\Omega.$
Lemma 3.5.
We assume $\tilde{A}\in[C^{0}(\overline{\Omega})]^{d\times d}$. Let $p\in(1,+\infty)$. For any $\delta>0$, there exists $R_{\delta}>0$ and $h_{\delta}>0$ such that for any $\boldsymbol{x}_{0}\in\overline{\Omega}$ with $\tilde{A}_{0}=\tilde{A}(\boldsymbol{x}_{0})$,
$\|(\mathcal{L}_{\tilde{A},h}-\mathcal{L}_{\tilde{A}_{0},h})w\|_{L_{h}^{p}(B_{R_{\delta}}(\boldsymbol{x}_{0}))}\leq\delta\|w\|_{W_{h}^{2,p}(B_{R_{\delta}}(\boldsymbol{x}_{0}))},\qquad\forall w\in W_{h}^{(p)},h\leq h_{\delta}.$
Proof.
We denote by $B_{t}:=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{0}|<t\}\subset\Omega$ for any $t>0$. For any $v_{h}\in V_{h}(B_{t})$, we have
$((\mathcal{L}_{\tilde{A},h}-\mathcal{L}_{\tilde{A}_{0},h})w,v_{h})_{\Omega}=\int_{\Omega}\left((\tilde{A}_{0}-\tilde{A}):\overline{\nabla}_{h}(\nabla w)\right)v_{h}d\boldsymbol{x}.$
We recall that $V_{h}(B_{t})=\{v\in V_{h}:v|_{\Omega\setminus B_{t}}=0\}$. We define $\mathcal{T}_{t,h}=\{K\in\mathcal{T}_{h}:\mathcal{H}^{d}(K\setminus B_{t})=0\}$, where $\mathcal{H}^{d}(\cdot)$ is the $d$-dimensional Hausdorff measure.
We define $p^{\prime}\in(1,+\infty)$ such that $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. So we have
$((\mathcal{L}_{\tilde{A},h}-\mathcal{L}_{\tilde{A}_{0},h})w,v_{h})_{\Omega}=\Sigma_{K\in\mathcal{T}_{t,h}}\int_{K}\left((\tilde{A}_{0}-\tilde{A}):\overline{\nabla}_{h}(\nabla w)\right)v_{h}d\boldsymbol{x}$
$\leq$ $\|\tilde{A}_{0}-\tilde{A}\|_{L^{\infty}(B_{t})}\left(\Sigma_{K\in\mathcal{T}_{t,h}}\|\overline{\nabla}_{h}(\nabla w)\|_{L^{p}(K)}^{p}\right)^{\frac{1}{p}}\|v_{h}\|_{L^{p^{\prime}}(\Omega)}.$
We denote by $\mathcal{F}_{t,h}$ the collection of all mesh interface of $\mathcal{T}_{t,h}$. By the construction of $\mathcal{T}_{t,h}$, we have
$\mathcal{H}^{d-1}(F\cap\overline{B_{t}})=\mathcal{H}^{d-1}(F),\qquad\forall F\in\mathcal{F}_{t,h}.$
Then with Definition 2.1 and (7), we have
$\Sigma_{K\in\mathcal{T}_{t,h}}\|\overline{\nabla}_{h}(\nabla w)\|_{L^{p}(K)}^{p}$
$\leq$ $C_{*}\left(\Sigma_{K\in\mathcal{T}_{t,h}}\|D_{h}^{2}w\|_{L^{p}(K)}^{p}+\Sigma_{F\in(\mathcal{F}_{t,h}\cap\mathcal{F}_{h}^{I})}h_{F}^{1-p}\|[\nabla w]\|_{L^{p}(F)}^{p}\right)\leq C_{*}\|w\|_{W_{h}^{2,p}(B_{t})}^{p},$
where $\mathcal{F}_{h}^{I}$ is the collection of all interior mesh faces of $\mathcal{T}_{h}$. The constant $C_{*}$ is independent of $\boldsymbol{x}_{0}$ and $t$.
We choose $t_{0}>0$ small enough, which is independent of $h$, such that
$\|\tilde{A}_{0}-\tilde{A}\|_{L^{\infty}(B_{t_{0}})}\leq\dfrac{\delta}{C_{*}+1}.$
Since $\tilde{A}\in[C(\overline{\Omega})]^{d\times d}$, $t_{0}$ can be chosen to be independent of $\boldsymbol{x}_{0}\in\overline{\Omega}$.
Then we have
$|((\mathcal{L}_{\tilde{A},h}-\mathcal{L}_{\tilde{A}_{0},h})w,v_{h})_{\Omega}|\leq\delta\|w\|_{W_{h}^{2,p}(B_{t_{0}})}^{p}\|v_{h}\|_{L^{p^{\prime}}(\Omega)},\qquad\forall v_{h}\in V_{h}(B_{t_{0}}).$
We choose $h_{\delta}=\frac{t_{0}}{4}$ such that for any $0<h\leq h_{\delta}$, $V_{h}(B_{t_{0}})$ is not empty. Then we can conclude that the proof is complete. ∎
Lemma 3.6.
We assume $\tilde{A}\in[C^{0}(\overline{\Omega})]^{d\times d}$ is uniformly elliptic. There exists $R_{1}>0$ and $h_{1}>0$ such that for any $\boldsymbol{x}_{0}\in\overline{\Omega}$,
(37) $\lambda_{0}\|w_{h}\|_{W_{h}^{2,p}(B_{R_{1}}(\boldsymbol{x}_{0}))}\leq\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(B_{2R_{1}}(\boldsymbol{x}_{0}))},\qquad\forall w_{h}\in V_{h}(B_{R_{1}}(\boldsymbol{x}_{0})),h<h_{1}.$
Here $p$ is valid in the range described in (1b). The constant $\lambda_{0}$ may depend on $\tilde{A}(\boldsymbol{x}_{0})$.
Proof.
The proof is an immediate consequence of Lemma 3.4 and Lemma 3.5. ∎
Lemma 3.7.
If $\tilde{A}\in[C^{0}(\overline{\Omega})]^{d\times d}$ is uniformly elliptic, then there is $h_{2}>0$ which may depend on $p$, such that for all $h\in(0,h_{2})$,
(38) $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\left(\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\Omega)}+\|w_{h}\|_{W^{1,p}(\Omega)}\right),\qquad\forall w_{h}\in V_{h}.$
Here $p$ is valid in the range described in (1b).
Proof.
We dived the proof into two steps.
Step $1$. For any $\boldsymbol{x}_{0}\in\overline{\Omega}$, we denote by $B_{R}:=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{0}|<R\}$ for any $R>0$. Let $\eta\in C^{3}(\mathbb{R}^{d})$ be a cut-off function satisfying
(39) $0\leq\eta\leq 1,\quad\eta|_{B_{\frac{4}{3}R}}=1,\quad\eta|_{\mathbb{R}^{d}\setminus B_{\frac{5}{3}R}}=0,\quad\|\eta\|_{W^{m,\infty}(\mathbb{R}^{d})}\leq CR^{-m}(m=0,1,2).$
From now on, we assume $h<\frac{R}{24}$. We define $\tilde{A}_{0}=\tilde{A}(\boldsymbol{x}_{0})\in\mathbb{R}^{d\times d}$.
We choose $w_{h}\in V_{h}$ arbitrarily. We notice that $\eta w_{h}\in W_{h}^{(p)}(B_{\frac{5}{3}R})$ and $I_{h}(\eta w_{h})\in V_{h}(B_{\frac{11}{6}R})=\{v_{h}\in V_{h}:v_{h}|_{\Omega\setminus B_{\frac{11}{6}R}}=0\}$. Here $I_{h}:C^{0}(\overline{\Omega})\rightarrow V_{h}$ is the standard nodal interpolant onto $V_{h}$.
We notice that $I_{h}(\eta v_{h})=\eta v_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$ for any $v_{h}\in V_{h}$. Thus [14, Lemma $2.4$] implies that for any $q\in(1,+\infty)$, there is a constant $C_{q}$ such that
(40) $\|I_{h}(\eta v_{h})-\eta v_{h}\|_{L^{q}(\Omega)}\leq C_{q}hR^{-1}\|v_{h}\|_{L^{q}(B_{2R})},$
$\|\nabla(I_{h}(\eta v_{h})-\eta v_{h})\|_{L^{q}(\Omega)}\leq C_{q}R^{-1}\|v_{h}\|_{L^{q}(B_{2R})},$
$\|I_{h}(\eta v_{h})-\eta v_{h}\|_{W_{h}^{2,q}(\Omega)}\leq C_{q}R^{-2}\|v_{h}\|_{W^{1,q}(B_{2R})},\quad\forall v_{h}\in V_{h}.$
By Lemma 3.6, there are two constants $\lambda_{0}>0$ and $R>0$ (we choose $R$ to be $R_{1}$ in Lemma 3.6) such that
(41) $\lambda_{0}\|w_{h}\|_{W_{h}^{p}(B_{R})}=\lambda_{0}\|I_{h}(\eta w_{h})\|_{W_{h}^{p}(B_{R})}\leq\|\mathcal{L}_{\tilde{A},h}(I_{h}(\eta w_{h}))\|_{L_{h}^{p}(B_{2R})}.$
Here the constant $\lambda_{0}$ may depend on $\tilde{A}(\boldsymbol{x}_{0})$.
By (41,8) and Definition 2.1, we have
$\lambda_{0}\|w_{h}\|_{W_{h}^{p}(B_{R})}=\lambda_{0}\|I_{h}(\eta w_{h})\|_{W_{h}^{p}(B_{R})}$
$\leq$ $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}+\|\mathcal{L}_{\tilde{A},h}(I_{h}(\eta w_{h})-\eta w_{h})\|_{L_{h}^{p}(B_{2R})}$
$\leq$ $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}+\|\mathcal{L}_{\tilde{A},h}(I_{h}(\eta w_{h})-\eta w_{h})\|_{L^{p}(\Omega)}$
$\leq$ $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}+\|\tilde{A}\|_{L^{\infty}(\Omega)}\|\overline{\nabla}_{h}\nabla(I_{h}(\eta w_{h})-\eta w_{h})\|_{L^{p}(\Omega)}$
$\leq$ $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}+C\|\tilde{A}\|_{L^{\infty}(\Omega)}\|I_{h}(\eta w_{h})-\eta w_{h}\|_{W_{h}^{2,p}(\Omega)}$
Then by (40), we have
(42) $\lambda_{0}\|w_{h}\|_{W_{h}^{p}(B_{R})}\leq\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}+CR^{-2}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
We define $p^{\prime}\in(1,+\infty)$ such that $\frac{1}{p}+\frac{1}{p^{\prime}}$. By (8),
$\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}=\sup_{v_{h}\in V_{h}(B_{2R})}\dfrac{(\mathcal{L}_{\tilde{A},h}(\eta w_{h}),v_{h})_{\Omega}}{\|v_{h}\|_{L^{p^{\prime}}(B_{2R})}},$
where $V_{h}(B_{2R})=\{v_{h}\in V_{h}:v_{h}|_{\Omega\setminus B_{2R}}=0\}$. We choose $v_{h}\in V_{h}(B_{2R})$ such that
(43) $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}=(\mathcal{L}_{\tilde{A},h}(\eta w_{h}),v_{h})_{\Omega},\quad\|v_{h}\|_{L^{p^{\prime}}(B_{2R})}=\|v_{h}\|_{L^{p^{\prime}}(\Omega)}=1.$
By (43,36) and Definition 2.1, we have
(44) $\|\mathcal{L}_{\tilde{A},h}(\eta w_{h})\|_{L_{h}^{p}(B_{2R})}=(\tilde{A}:\overline{\nabla}_{h}(\nabla(\eta w_{h})),v_{h})_{\Omega}$
$=$ $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{\Omega}+(\tilde{A}:\overline{\nabla}_{h}(w_{h}\nabla\eta),v_{h})_{\Omega}$
$\leq$ $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{\Omega}+\|\tilde{A}\|_{L^{\infty}(\Omega)}\|\overline{\nabla}_{h}(w_{h}\nabla\eta)\|_{L^{p}(\Omega)}$
$\leq$ $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{\Omega}+C\|w_{h}\nabla\eta\|_{W_{h}^{2,p}(\Omega)}$
$\leq$ $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{B_{2R}}+CR^{-2}\|w_{h}\|_{W^{1,p}(\Omega)}.$
Therefore, (42,44) imply that
(45) $\lambda_{0}\|w_{h}\|_{W_{h}^{p}(B_{R})}\leq$ $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{\Omega}+CR^{-2}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
We denote by $\overline{\eta}_{h}\in P_{0}(\mathcal{T}_{h})$ the element-wise average of $\eta$. Then we have
(46) $(\tilde{A}:\overline{\nabla}_{h}(\eta\nabla w_{h}),v_{h})_{\Omega}$
$=$ $(\tilde{A}:\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h}),v_{h})_{\Omega}+(\tilde{A}:\left(\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\right),v_{h})_{\Omega}$
$\qquad+(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h}))_{\Omega}+(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),I_{h}(\eta v_{h}))_{\Omega}$
We give estimates of all four terms on the right hand side of (46) in the following.
By (43), we have
$(\tilde{A}:\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h}),v_{h})_{\Omega}\leq\|\tilde{A}\|_{L^{\infty}(\Omega)}\|\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h})\|_{L^{p}(\Omega)}$
Since $\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h})\in[\overline{V}_{h}]^{d\times d}$, we denote by $\underline{\varphi}_{h}\in[\overline{V}_{h}]^{d\times d}$ such that
$\|\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h})\|_{L^{p}(\Omega)}=\left(\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h}),\underline{\varphi}_{h}\right)_{\Omega},\qquad\|\underline{\varphi}_{h}\|_{L^{p^{\prime}}(\Omega)}=1.$
By Definition 2.1 and the fact $\eta=\overline{\eta}_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$, we have
$\left(\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h}),\underline{\varphi}_{h}\right)_{\Omega}=-\langle[(\eta-\overline{\eta}_{h})\nabla w_{h}]\otimes\boldsymbol{n},\{\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}+(\nabla((\eta-\overline{\eta}_{h})\nabla w_{h}),\underline{\varphi}_{h})_{\mathcal{T}_{h}}$
$=$ $-\langle[\eta-\overline{\eta}_{h}]\{\nabla w_{h}\}\otimes\boldsymbol{n},\{\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}-\langle\{\eta-\overline{\eta}_{h}\}[\nabla w_{h}]\otimes\boldsymbol{n},\{\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}+(\nabla((\eta-\overline{\eta}_{h})\nabla w_{h}),\underline{\varphi}_{h})_{\mathcal{T}_{h}}$
$\leq$ $C\left(R^{-1}\|w_{h}\|_{W^{1,p}(B_{\frac{11}{6}R})}+hR^{-1}\|w_{h}\|_{W_{h}^{2,p}(B_{\frac{11}{6}R})}\right)\|\underline{\varphi}_{h}\|_{L^{p^{\prime}}(\Omega)}.$
By [14, Lemma $2.2$], we have
$\|_{W_{h}^{2,p}(B_{\frac{11}{6}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
Therefore, we have
(47) $(\tilde{A}:\overline{\nabla}_{h}((\eta-\overline{\eta}_{h})\nabla w_{h}),v_{h})_{\Omega}\leq CR^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
By (43), we have
$(\tilde{A}:\left(\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\right),v_{h})_{\Omega}\leq\|\tilde{A}\|_{L^{\infty}(\Omega)}\|\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(\Omega)}.$
We notice $\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\in[\overline{V}_{h}]^{d\times d}$. we denote by $\underline{\varphi}_{h}\in[\overline{V}_{h}]^{d\times d}$ such that
$\|\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(\Omega)}=\left(\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h}),\underline{\varphi}_{h}\right)_{\Omega},\qquad\|\underline{\varphi}_{h}\|_{L^{p^{\prime}}(\Omega)}=1.$
Since $\overline{\eta}_{h}\underline{\varphi}_{h}\in[\overline{V}_{h}]^{d\times d}$, Definition 2.1 implies that
$\left(\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h}),\underline{\varphi}_{h}\right)_{\Omega}$
$=$ $-\langle[\overline{\eta}_{h}\nabla w_{h}]\otimes\boldsymbol{n},\{\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}+\langle[\nabla w_{h}]\otimes\boldsymbol{n},\{\overline{\eta}_{h}\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}$
$=$ $-\langle[\overline{\eta}_{h}]\{\nabla w_{h}\}\otimes\boldsymbol{n},\{\underline{\varphi}_{h}\}\rangle_{\mathcal{F}_{h}^{I}}+\frac{1}{4}\langle[\nabla w_{h}]\otimes\boldsymbol{n},[\overline{\eta}_{h}][\underline{\varphi}_{h}]\rangle_{\mathcal{F}_{h}^{I}}.$
Then by discrete trace inequality and the fact $\eta=\overline{\eta}_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$, we have
(48) $(\tilde{A}:\left(\overline{\nabla}_{h}(\overline{\eta}_{h}\nabla w_{h})-\overline{\eta}_{h}\overline{\nabla}_{h}(\nabla w_{h})\right),v_{h})_{\Omega}\leq CR^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
By the fact $\eta=\overline{\eta}_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$ and Definition 2.1,
$(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h}))_{\Omega}$
$\leq$ $\|\tilde{A}\|_{L^{\infty}(\Omega)}\|\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(B_{\frac{11}{6}R})}\|\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h})\|_{L^{p^{\prime}}(\Omega)}$
$\leq$ $C\|w_{h}\|_{W_{h}^{2,p}(B_{\frac{23}{12}}R)}\|\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h})\|_{L^{p^{\prime}}(\Omega)}.$
By [14, Lemma $2.2$], we have
$\|_{W_{h}^{2,p}(B_{\frac{23}{12}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
So we have
$(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h}))_{\Omega}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}\|\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h})\|_{L^{p^{\prime}}(\Omega)}.$
By (40,43), we have
(49) $(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),\overline{\eta}_{h}v_{h}-I_{h}(\eta v_{h}))_{\Omega}$
$\leq$ $CR^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}\|v_{h}\|_{L^{p^{\prime}}(\Omega)}=CR^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
We notice $I_{h}(\eta v_{h})\in V_{h}(B_{\frac{11}{6}R})$. Then by (36,8,40,43), we have
(50) $(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),I_{h}(\eta v_{h}))_{\Omega}=(\mathcal{L}_{\tilde{A},h}w_{h},I_{h}(\eta v_{h}))_{\Omega}$
$\leq$ $\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(B_{2R})}\|I_{h}(\eta v_{h})\|_{L^{p^{\prime}}(\Omega)}\leq C\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(B_{2R})}\|v_{h}\|_{L^{p^{\prime}}(B_{2R})}=C\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(B_{2R})}.$
According to (45,46, 47,48,49, 50), we have
(51) $\lambda_{0}\|w_{h}\|_{W_{h}^{2,p}(B_{R})}\leq C\left(\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(B_{2R})}+R^{-2}\|w_{h}\|_{W^{1,p}(B_{2R})}\right),\qquad\forall w_{h}\in V_{h}.$
We recall that $R$ is chosen to be $R_{1}$ in Lemma 3.6. The constant $\lambda_{0}$ may depend on $\tilde{A}(\boldsymbol{x}_{0})$.
Step $2$. Let $\{\boldsymbol{x}_{j}\}_{j=1}^{N}\subset\overline{\Omega}$ such that $\overline{\Omega}=\cup_{j=1}^{N}\overline{B}_{R_{j}^{\prime}}(\boldsymbol{x}_{j})$ where $R_{j}^{\prime}$ is the constant $R$ for $\boldsymbol{x}_{j}$ in (51), and $B_{R_{j}^{\prime}}:=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{j}|<R_{j}^{\prime}\}$. Since $\overline{\Omega}$ is compact, $N$ is finite and independent of $h$.
We denote by $S_{j}=B_{R_{j}^{\prime}}$ and $\tilde{S}_{j}=B_{2R_{j}^{\prime}}:=\{\boldsymbol{x}\in\Omega:|\boldsymbol{x}-\boldsymbol{x}_{j}|<2R_{j}^{\prime}\}$ for any $1\leq j\leq N$.
Since $R_{j}$ is independent of $h$ for any $1\leq j\leq N$, (51) implies that
$\|w_{h}\|_{W_{h}^{2,p}(\Omega)}^{p}\leq\Sigma_{j=1}^{N}\|w_{h}\|_{W_{h}^{2,p}(S_{j})}^{p}\leq C\left(\Sigma_{j=1}^{N}\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\tilde{S}_{j})}^{p}+\|w_{h}\|_{W^{1,p}(\Omega)}^{p}\right).$
By (8),
$\Sigma_{j=1}^{N}\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\tilde{S}_{j})}^{p}=\Sigma_{j=1}^{N}\left|\sup_{v_{h}\in V_{h}(\tilde{S}_{j})}\dfrac{(\mathcal{L}_{\tilde{A},h}w_{h},v_{h})_{\Omega}}{\|v_{h}\|_{L^{p^{\prime}}(\tilde{S}_{j})}}\right|^{p}$
$\leq$ $N\left|\sup_{v_{h}\in V_{h}}\dfrac{(\mathcal{L}_{\tilde{A},h}w_{h},v_{h})_{\Omega}}{\|v_{h}\|_{L^{p^{\prime}}(\Omega)}}\right|^{p}=N\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\Omega)}.$
Therefore we can conclude that the proof is complete. ∎
Lemma 3.8.
(Discrete compactness) Let $1<p<+\infty$ and $\Omega$ be a Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). We assume that there are $0<M_{1}<+\infty$ and $\{w_{h}\in V_{h}\}_{h>0}$ satisfying
$\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1},\qquad\forall h>0.$
Then there is $w\in W^{2}\cap W_{0}^{1,p}(\Omega)$ and $0<M_{2}<+\infty$ such that
$\|w_{h}-w\|_{W^{1,p}(\Omega)}\rightarrow 0,\qquad\overline{\nabla}_{h}(\nabla w_{h})\rightharpoonup D^{2}w\text{ in }[L^{p}(\Omega)]^{d\times d},$
for a subsequence of $h\rightarrow 0$; and
$\quad\|\overline{\nabla}_{h}(\nabla w)\|_{L^{p}(\Omega)}+\|\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(\Omega)}\leq M_{2},\qquad\forall h>0.$
Proof.
Since
$\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1},\qquad\forall h>0,$
there is $w\in W_{0}^{1,p}(\Omega)$ such that
(52) $\|w_{h}-w\|_{L^{p}(\Omega)}\rightarrow 0,\qquad w_{h}\rightharpoonup w\text{ in }W_{0}^{1,p}(\Omega),$
for a subsequence of $h\rightarrow 0$. We denote by $\boldsymbol{\phi}_{h}\in[H^{1}(\Omega)\cap\overline{V}_{h}]^{d}$ the standard $H^{1}$-conforming interpolation of $\nabla w_{h}$. By Lemma 2.3, we have
(53) $\|\boldsymbol{\phi}_{h}-\nabla w_{h}\|_{L^{p}(\Omega)}\leq Ch\left(\Sigma_{F\in\mathcal{F}_{h}^{I}}h_{F}^{1-p}\|[\nabla w_{h}]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}.$
(53) and the fact $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1}$ imply that there is $\boldsymbol{\phi}\in[W^{1,p}(\Omega)]^{d}$ such that
(54) $\|\boldsymbol{\phi}_{h}-\boldsymbol{\phi}\|_{L^{p}(\Omega)}+\|\nabla w_{h}-\boldsymbol{\phi}\|_{L^{p}(\Omega)}\rightarrow 0,\quad\text{ and }\quad\boldsymbol{\phi}_{h}\rightharpoonup\boldsymbol{\phi}\text{ in }[W^{1,p}(\Omega)]^{d},$
for a subsequence of $h\rightarrow 0$. (52, 54) imply that $\nabla w=\boldsymbol{\phi}$ almost everywhere in $\Omega$. Therefore, we have that $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ and
(55) $\|w_{h}-w\|_{W^{1,p}(\Omega)}+\|\boldsymbol{\phi}_{h}-\nabla w\|_{L^{p}(\Omega)}\rightarrow 0,\quad\text{ and }\boldsymbol{\phi}_{h}\rightharpoonup\nabla w\text{ in }[W^{1,p}(\Omega)]^{d},$
for a subsequence of $h\rightarrow 0$.
According to Definition 2.1 with the fact $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1}<+\infty$ and $w\in W^{2,p}(\Omega)$, there is a constant $M_{2}$ such that
(56) $\|\overline{\nabla}_{h}(\nabla w)\|_{L^{p}(\Omega)}+\|\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(\Omega)}\leq M_{2}<+\infty.$
We denote by $P_{h}$ the standard $L^{2}$-orthogonal projection onto $V_{h}$. By Lemma 2.2, there is a constant $\tilde{C}_{1}$ independent of $h$ such that for any $v\in W_{0}^{1,p^{\prime}}(\Omega)$,
(57) $\|P_{h}v\|_{L^{p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{L^{p^{\prime}}(\Omega)},\qquad\|P_{h}v\|_{W^{1,p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{W^{1,p^{\prime}}(\Omega)}.$
By (56), there is $\underline{\psi}\in[L^{p}(\Omega)]^{d\times d}$ such that
$\overline{\nabla}_{h}(\nabla w_{h})\rightharpoonup\underline{\psi}\in[L^{p}(\Omega)]^{d\times d},$
for a subsequence of $h\rightarrow 0$. On the other hand, we have that for any $\underline{\zeta}\in[C_{0}^{\infty}(\Omega)]^{d\times d}$,
$(\overline{\nabla}_{h}(\nabla w_{h}),\underline{\zeta})_{\Omega}=(\overline{\nabla}_{h}(\nabla w_{h}),P_{h}(\underline{\zeta}))_{\Omega}+(\overline{\nabla}_{h}(\nabla w_{h}),P_{h}(\underline{\zeta})-\underline{\zeta})_{\Omega}.$
Since $P_{h}(\underline{\zeta})\in[V_{h}]^{d\times d}$, then by Definition 2.1
$(\overline{\nabla}_{h}(\nabla w_{h}),\underline{\zeta})_{\Omega}=-(\nabla w_{h},\nabla\cdot P_{h}(\underline{\zeta}))_{\Omega}+(\overline{\nabla}_{h}(\nabla w_{h}),P_{h}(\underline{\zeta})-\underline{\zeta})_{\Omega}$
$=$ $-(\nabla w_{h},\nabla\cdot\underline{\zeta})_{\Omega}-(\nabla w_{h},\nabla\cdot(P_{h}(\underline{\zeta})-\underline{\zeta}))_{\Omega}+(\overline{\nabla}_{h}(\nabla w_{h}),P_{h}(\underline{\zeta})-\underline{\zeta})_{\Omega}$
In the following of the proof, $\lim_{h\rightarrow 0}$ means the limit when a subsequence mentioned above of $h\rightarrow 0$. By (55), we have
$\lim_{h\rightarrow 0}-(\nabla w_{h},\nabla\cdot\underline{\zeta})_{\Omega}=-(\nabla w,\nabla\cdot\underline{\zeta})_{\Omega}=(D^{2}w,\underline{\zeta})_{\Omega}.$
By (56) and the fact $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1}<+\infty$, (57) implies that
$\lim_{h\rightarrow 0}(\nabla w_{h},\nabla\cdot(P_{h}(\underline{\zeta})-\underline{\zeta}))_{\Omega}=\lim_{h\rightarrow 0}(\overline{\nabla}_{h}(\nabla w_{h}),P_{h}(\underline{\zeta})-\underline{\zeta})_{\Omega}=0.$
Therefore, we have
$\lim_{h\rightarrow 0}(\overline{\nabla}_{h}(\nabla w_{h}),\underline{\zeta})_{\Omega}=(D^{2}w,\underline{\zeta})_{\Omega},\qquad\forall\underline{\zeta}\in[C_{0}^{\infty}(\Omega)]^{d\times d}.$
Therefore, we have $\underline{\psi}=D^{2}w$ almost everywhere in $\Omega$. So we have
(58) $\overline{\nabla}_{h}(\nabla w_{h})\rightharpoonup D^{2}w\in[L^{p}(\Omega)]^{d\times d},$
for a subsequence of $h\rightarrow 0$.
According to (55, 56, 58), We can conclude the proof is complete. ∎
Theorem 3.3.
(Stability of the FEM (14)) Let $A\in[C^{0}(\overline{\Omega})]^{d\times d}$ uniformly elliptic, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ is a bounded open Lipschitz polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). There is $h_{2}>0$ which may depend on $p$, such that for all $h\in(0,h_{2})$,
(59) $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)},\qquad\forall w_{h}\in V_{h}.$
Here $p$ is valid in the range described in (1b). $\tilde{A}$, $\tilde{\boldsymbol{b}}$ and $\tilde{c}$ are defined in (13). The norm $\|\cdot\|_{L_{h}^{p}(\Omega)}$ is defined in (8).
Proof.
We notice that by (12), $A\in[C(\overline{\Omega})]^{d\times d}$ is equivalent to $\tilde{A}\in[C(\overline{\Omega})]^{d\times d}$, and $c\leq 0$ is equivalent to $\tilde{c}\leq 0$.
For the sake of simplicity, we only give detailed proof of the case $d=3$. The proof of the case $d=2$ ($\Omega$ can be non-convex) can be proven in the same manner.
According to Lemma 3.7, it is sufficient to prove that for any $p\in(1,2]$,
$\|w_{h}\|_{W^{1,p}(\Omega)}\leq C\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)},\qquad\forall w_{h}\in V_{h}.$
Here the constant $C$ may depend on $p$.
We prove by contradiction. Without losing generality, we assume that there is one $p\in(1,2]$ and a sequence of $\{w_{h}\in V_{h}\}_{h>0}$, such that
(60) $\|w_{h}\|_{W^{1,p}(\Omega)}=1,\quad\text{and }\quad\lim_{h\rightarrow 0}\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)}=0.$
So there is a positive constant $C$ such that
$\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\Omega)}=\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})\|_{L_{h}^{p}(\Omega)}$
$\leq$ $C\left(\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)}+\|w_{h}\|_{W^{1,p}(\Omega)}\right).$
The equality above is due to (36). According to Lemma 3.7, there is $M_{1}<+\infty$ such that
$\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1},\qquad\forall h>0.$
Therefore, Lemma 3.8 implies that there is $w\in W^{2}\cap W_{0}^{1,p}(\Omega)$ and $0<M_{2}<+\infty$ such that
(61) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\rightarrow 0,$
(62) $\overline{\nabla}_{h}(\nabla w_{h})\rightharpoonup D^{2}w\text{ in }L^{p}(\Omega),$
for a subsequence of $h\rightarrow 0$; and
(63) $\quad\|\overline{\nabla}_{h}(\nabla w)\|_{L^{p}(\Omega)}+\|\overline{\nabla}_{h}(\nabla w_{h})\|_{L^{p}(\Omega)}\leq M_{2},\qquad\forall h>0.$
In the following of the proof, we don’t distinguish $h\rightarrow 0$ and a subsequence of $h\rightarrow 0$. Then (60) and (61) imply that $\|w\|_{W^{1,p}(\Omega)}\geq\frac{1}{2}$. By Theorem 3.2, in order to achieve contradiction with $\|w\|_{W^{1,p}(\Omega)}\geq\frac{1}{2}$, we only need to prove that
$\tilde{A}:D^{2}w+\tilde{\boldsymbol{b}}\cdot\nabla w+\tilde{c}w=0\text{ almost everywhere in }\Omega.$
Since $w\in W^{2,p}(\Omega)$, it is sufficient to prove that
(64) $(\tilde{A}:D^{2}w+\tilde{\boldsymbol{b}}\cdot\nabla w+\tilde{c}w,v)_{\Omega}=0,\qquad\forall v\in C_{0}^{\infty}(\Omega).$
We choose $v\in C_{0}^{\infty}(\Omega)$ arbitrarily. It is easy to see
$(\tilde{A}:D^{2}w+\tilde{\boldsymbol{b}}\cdot\nabla w+\tilde{c}w,v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}w-\overline{\nabla}_{h}(\nabla w_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}),v)_{\Omega}+(\tilde{\boldsymbol{b}}\cdot\nabla w+\tilde{c}w,v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}w-\overline{\nabla}_{h}(\nabla w_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h},P_{h}v)_{\Omega}$
$\qquad+(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h},v-P_{h}v)_{\Omega}+(\tilde{\boldsymbol{b}}\cdot\nabla(w-w_{h})+\tilde{c}(w-w_{h}),v)_{\Omega}.$
Here we denote by $P_{h}$ the standard $L^{2}$-orthogonal projection onto $V_{h}$. We define $p^{\prime}\in(1,+\infty)$ by $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. By (62) and the fact that $\tilde{A}v\in[L^{p^{\prime}}(\Omega)]^{d\times d}$,
$\lim_{h\rightarrow 0}(\tilde{A}:(D^{2}w-\overline{\nabla}_{h}(\nabla w_{h})),v)_{\Omega}=0.$
By Lemma 2.2, there is a constant $\tilde{C}_{1}$ independent of $h$,
(65) $\|P_{h}v\|_{L^{p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{L^{p^{\prime}}(\Omega)},\qquad\|P_{h}v\|_{W^{1,p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{W^{1,p^{\prime}}(\Omega)}.$
Therefore, by (60) and (65), we have
$\lim_{h\rightarrow 0}|(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h}+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}),P_{h}v)_{\Omega}|$
$\leq$ $\lim_{h\rightarrow 0}\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)}\|P_{h}v\|_{L^{p^{\prime}}(\Omega)}$
$\leq$ $\tilde{C}_{1}\lim_{h\rightarrow 0}\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)}\|v\|_{L^{p^{\prime}}(\Omega)}=0.$
By (60,63,65), we have
$\lim_{h\rightarrow 0}(\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h},v-P_{h}v)_{\Omega}=0.$
By (61), we have
$\lim_{h\rightarrow 0}(\tilde{\boldsymbol{b}}\cdot\nabla(w-w_{h})+\tilde{c}(w-w_{h}),v)_{\Omega}=0.$
Therefore, we have
(66) $(\tilde{A}:D^{2}w+\tilde{\boldsymbol{b}}\cdot\nabla w+\tilde{c}w,v)_{\Omega}=0,\qquad\forall v\in C_{0}^{\infty}(\Omega).$
(66) implies that the proof is complete. ∎
3.3. Well-posedness of linear elliptic PDE in non-divergence form with uniformly continuous $A$ and optimal convergence of the numerical solution of FEM
Theorem 3.4.
Let $A\in[C^{0}(\overline{\Omega})]^{d\times d}$ uniformly elliptic, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ is a bounded open Lipschitz polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). For any $f\in L^{p}(\Omega)$, there is a unique $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ such that
(67a) $A:D^{2}u+\boldsymbol{b}\cdot\nabla u+cu=f\text{ in }\Omega,$
(67b) $\|u\|_{W^{2,p}(\Omega)}\leq C_{p}\|f\|_{L^{p}(\Omega)}.$
Here the constant $C_{p}$ is the same as the one in Theorem 3.2. $p$ is valid in the range described in (1a).
Let $u_{h}\in V_{h}$ be the numerical solution of the finite element method (14). Then we have
(68) $\|u_{h}-u\|_{W^{1,p}(\Omega)}\rightarrow 0,\qquad\overline{\nabla}_{h}(\nabla u_{h})\rightharpoonup D^{2}u\text{ in }[L^{p}(\Omega)]^{d\times d},$
as $h\rightarrow 0$. And if $h\in(0,h_{2})$ where $h_{2}$ is introduced in Theorem 3.3, then for any $\chi_{h}\in V_{h}$,
(69) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big).$
Here $p$ is valid in the range described in (1b). $\overline{P}_{h}$ denotes the standard $L^{2}$-orthogonal projection onto $\overline{V}_{h}]^{d\times d}$.
Proof.
We divide the proof into two steps. In step $1$, We will first prove the estimates (67) and (68,69) for $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$. In step $2$, we will prove the estimate (67) for $1<p\leq\frac{4}{3}-\epsilon_{1}$ and $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
Step $1$. We consider $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
By the design of the finite element method (14) and Theorem 3.3, there is a unique $u_{h}\in V_{h}$ to be the numerical solution of (14) and
$\|u_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|f\|_{L^{p}(\Omega)},$
if $h$ is small enough. Then by Lemma 3.8, there is $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ such that
$\|u_{h}-u\|_{W^{1,p}(\Omega)}\rightarrow 0,\qquad\overline{\nabla}_{h}(\nabla u_{h})\rightharpoonup D^{2}u\text{ in }[L^{p}(\Omega)]^{d\times d},$
for a sequence of $h\rightarrow 0$; and there is $0<M_{2}<+\infty$ such that $\|\overline{\nabla}(\nabla u_{h})\|_{L^{p}(\Omega)}\leq M_{2}$ for any $h>0$.
Now we need to prove (67a). We define $p^{\prime}\in(1,+\infty)$ by $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. We denote by $P_{h}$ the standard $L^{2}$-orthogonal projection onto $V_{h}$. It is well-known that there is a constant $\tilde{C}_{1}$ independent of $h$,
(70) $\|P_{h}v\|_{L^{p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{L^{p^{\prime}}(\Omega)},\qquad\|P_{h}v\|_{W^{1,p^{\prime}}(\Omega)}\leq\tilde{C}_{1}\|v\|_{W^{1,p^{\prime}}(\Omega)}.$
We choose $v\in C_{0}^{\infty}(\Omega)$ arbitrarily. By the design of the finite element method (14), We have
$(\tilde{A}:D^{2}u+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}u-\overline{\nabla}(\nabla u_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}u-\overline{\nabla}(\nabla u_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v-P_{h}v)_{\Omega}$
$\qquad+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,P_{h}v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}u-\overline{\nabla}(\nabla u_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v-P_{h}v)_{\Omega}$
$\qquad+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u_{h}+\tilde{c}u_{h},P_{h}v)_{\Omega}+(\tilde{\boldsymbol{b}}\cdot\nabla(u-u_{h})+\tilde{c}(u-u_{h}),P_{h}v)_{\Omega}$
$=$ $(\tilde{A}:(D^{2}u-\overline{\nabla}(\nabla u_{h})),v)_{\Omega}+(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v-P_{h}v)_{\Omega}$
$\qquad+(\tilde{f},v)_{\Omega}+(\tilde{f},P_{h}v-v)_{\Omega}+(\tilde{\boldsymbol{b}}\cdot\nabla(u-u_{h})+\tilde{c}(u-u_{h}),P_{h}v)_{\Omega}.$
Since $\overline{\nabla}_{h}(\nabla u_{h})\rightharpoonup D^{2}u\text{ in }[L^{p}(\Omega)]^{d\times d}$, we have
$(\tilde{A}:(D^{2}u-\overline{\nabla}_{h}(\nabla u_{h})),v)_{\Omega}+(\tilde{\boldsymbol{b}}\cdot\nabla(u-u_{h})+\tilde{c}(u-u_{h}),P_{h}v)_{\Omega}\rightarrow 0,$
for a sequence of $h\rightarrow 0$.
By (70) and the fact that $\|\overline{\nabla}(\nabla u_{h})\|_{L^{p}(\Omega)}\leq M_{2}$, we have
$|(\tilde{A}:\overline{\nabla}(\nabla u_{h})+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v-P_{h}v)_{\Omega}|$
$\leq$ $C\left(\|\overline{\nabla}(\nabla u_{h})\|_{L^{p}(\Omega)}+\|u\|_{W^{1,p}(\Omega)}\right)\|v-P_{h}v\|_{L^{p^{\prime}}(\Omega)}\rightarrow 0,$
as $h\rightarrow 0$.
Similarly, we have
$|(\tilde{f},P_{h}v-v)_{\Omega}|\rightarrow 0,$
as $h\rightarrow 0$.
So we have
$(\tilde{A}:D^{2}u+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u,v)_{\Omega}=(\tilde{f},v)_{\Omega}.$
Since $v\in C_{0}^{\infty}(\Omega)$ is chosen arbitrarily, $\tilde{A}:D^{2}+\tilde{\boldsymbol{b}}\cdot\nabla u+\tilde{c}u=\tilde{f}$ in $\Omega$ almost everywhere. By the definition of the function $\gamma$ (see (12)) and the fact that $A$ is uniformly elliptic and uniformly bounded in $\Omega$, $\gamma$ is uniformly bounded from below and above by positive constants. So we have $A:D^{2}u+\boldsymbol{b}\cdot\nabla u+cu=f$ in $\Omega$ almost everywhere. So (67a) is proven. Theorem 3.2 implies the uniqueness of $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$, which results in
$\|u_{h}-u\|_{W^{1,p}(\Omega)}\rightarrow 0,\qquad\overline{\nabla}_{h}(\nabla u_{h})\rightharpoonup D^{2}u\text{ in }[L^{p}(\Omega)]^{d\times d},$
as $h\rightarrow 0$.
The proof of (69) is straightforward due to Theorem 3.3 and the construction of the finite element method (14). Thus the proof in Step $1$ is complete.
Step $2$. We consider $1<p\leq\frac{4}{3}-\epsilon_{1}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
Let $\{{f_{i}}\}_{i\geq 1}\subset L^{\frac{4}{3}}(\Omega)$ such that $\|f_{i}-f\|_{L^{p}(\Omega)}\rightarrow 0$ as $i\rightarrow+\infty$. For any $i\geq 1$, we denote by $u_{i}\in W^{2,\frac{4}{3}}(\Omega)\cap W_{0}^{1,\frac{4}{3}}(\Omega)$ satisfying
$A:D^{2}u_{i}+\boldsymbol{b}\cdot\nabla u_{i}+cu_{i}=f_{i}\text{ in }\Omega\text{ almost everywhere}.$
Since $\|f_{i}-f\|_{L^{p}(\Omega)}\rightarrow 0$ as $i\rightarrow+\infty$, $\{\|f_{i}\|_{L^{p}(\Omega)}\}_{i\geq 1}$ have a uniform upper bound. So $\{\|u_{i}\|_{W^{2,p}(\Omega)}\}_{i\geq 1}$ have a uniform upper bound as well. So there is $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ such that
$u_{i}\rightharpoonup u\text{ in }W^{2,p}(\Omega),$
for a subsequence of $i\rightarrow+\infty$.
We choose $v\in C_{0}^{\infty}(\Omega)$ arbitrarily. Then
$(A:D^{2}u+\boldsymbol{b}\cdot\nabla u+cu,v)_{\Omega}$
$=$ $(A:D^{2}(u-u_{i})+\boldsymbol{b}\cdot\nabla(u-u_{i})+c(u-u_{i}),v)_{\Omega}+(A:D^{2}u_{i}+\boldsymbol{b}\cdot\nabla u_{i}+cu_{i},v)_{\Omega}$
$=$ $(A:D^{2}(u-u_{i})+\boldsymbol{b}\cdot\nabla(u-u_{i})+c(u-u_{i}),v)_{\Omega}+(f_{i}-f,v)_{\Omega}+(f,v)_{\Omega}.$
Since $u_{i}\rightharpoonup u\text{ in }W^{2,p}(\Omega)$ for a subsequence of $i\rightarrow+\infty$ and $\|f_{i}-f\|_{L^{p}(\Omega)}\rightarrow 0$ as $i\rightarrow+\infty$, we have
$(A:D^{2}u+\boldsymbol{b}\cdot\nabla u+cu,v)_{\Omega}=(f,v)_{\Omega}.$
Since $v\in C_{0}^{\infty}(\Omega)$ is chosen arbitrarily, $A:D^{2}u=f$ in $\Omega$ almost everywhere. By Theorem 3.2, $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ is the unique solution of (2). Thus the proof of Step $2$ is complete. ∎
3.4. Linear elliptic PDE in non-divergence form with discontinuous $A$ but $\gamma A$ dominated by $I_{d}$
We skip the proof of Theorem 3.5 and Theorem 3.6, since they are special examples of Theorem 4.1 and Theorem 4.3 with the index set $\Lambda$ containing a single element.
Theorem 3.5.
Let $A\in[L^{\infty}(\Omega)]^{d\times d}$ uniformly elliptic, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ be a Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). There is a constant $0\leq\underline{\kappa}<1$ which may depend on $p$, such that if
(71) $\dfrac{|A(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}(\boldsymbol{x})|^{2}+\lambda^{-2}|c(\boldsymbol{x})|^{2}}{\left(\text{Tr}A(\boldsymbol{x})+\lambda^{-1}|c(\boldsymbol{x})|\right)^{2}}$
$\leq$ $\dfrac{1}{d+\epsilon},\qquad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},$
for some $\epsilon\in[\underline{\kappa},1)$ and $\lambda>0$, then there is a unique solution $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfying (2) for any $f\in L^{p}(\Omega)$. Furthermore, there is a positive constant $C$ which may depend on $p$, such that
$\|u\|_{W^{2,p}(\Omega)}\leq C\|f\|_{L^{p}(\Omega)}.$
Here $p$ is valid in the range described in (1a). In addition, $\underline{\kappa}=0$ if $p=2$ and the domain is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$).
Theorem 3.6.
Let $A\in[L^{\infty}(\Omega)]^{d\times d}$ uniformly elliptic, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, $c\in L^{\infty}(\Omega)$ with $c\leq 0$, and $\Omega$ be a Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). There is a constant $0\leq\kappa<1$ which my depend on $p$, such that if
(72) $\dfrac{|A(\boldsymbol{x})|^{2}(\text{Tr}A(\boldsymbol{x}))^{-2}}{1+2\lambda^{-1}|c(\boldsymbol{x})|(\text{Tr}A(\boldsymbol{x}))^{-1}-|A(\boldsymbol{x})|^{-2}\big(\lambda^{-2}|c(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}(\boldsymbol{x})|^{2}\big)}$
$\leq$ $\dfrac{1}{d+\epsilon},\quad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},$
for some constants $\epsilon\in[\kappa,1)$ and $\lambda>0$, then there is a unique numerical solution $u_{h}\in V_{h}$ of the finite element method (14) for any $0<h<h_{3}$. In addition, there is a positive constant $C$ which may depend on $p$, such that
$\|u_{h}\|_{W^{2,p}_{h}(\Omega)}\leq C\|f\|_{L^{p}(\Omega)}.$
Furthermore, if there are $\epsilon\in[0,1)$ and $\lambda>0$ such that (72) holds, then (71) in Theorem 3.5 holds as well for the same $\epsilon$ and $\lambda$. We also have the convergent result that for any $\chi_{h}\in V_{h}$,
(73) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
if $0<h<h_{3}$. Here $\overline{P}_{h}$ is the standard $L^{2}$-orthogonal projection onto $\overline{V}_{h}]^{d\times d}$. $p$ is valid in the range described in (1b).
4. Hamilton-Jacobi-Bellman equation
For the elliptic HJB equation (3), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (3), while Theorem 4.3 provides optimal convergence of the FEM (16).
The identity (74) in the following Lemma 4.1 is essentially important in analysis in Section 4, since it implies the index set $\Lambda$ can be replaced by a countably subset $\tilde{\Lambda}$.
Lemma 4.1.
We assume that the assumption (6) is satisfied. Then for any $B\in[L^{1}(\Omega)]^{d\times d},\boldsymbol{c}\in[L^{1}(\Omega)]^{d},w\in L^{1}(\Omega)$,
$\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]$
is a measurable function on $\Omega$, and
(74) $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]$
$=$ $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}[A^{\tilde{\alpha}}:B+\boldsymbol{b}^{\tilde{\alpha}}\cdot\boldsymbol{c}+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}]\qquad\text{ in }\Omega\text{ almost everywhere}.$
In addition, if (4,5) hold and $B\in[L^{p}(\Omega)]^{d\times d},\boldsymbol{c}\in[L^{p}(\Omega)]^{d},w\in L^{p}(\Omega)$ and $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ for some $p\in[1,+\infty)$, then we have that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and a positive constant $C$ satisfying
(75) $\|\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\|_{L^{p}(\Omega)}$
$\leq$ $C\big((\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\|A^{\tilde{\alpha}}\|_{L^{\infty}(\Omega)}\|B\|_{L^{p}(\Omega)}+\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\|\boldsymbol{b}^{\tilde{\alpha}}\|_{L^{\infty}(\Omega)}\|\boldsymbol{c}\|_{L^{p}(\Omega)}+\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\|c^{\tilde{\alpha}}\|_{L^{\infty}(\Omega)}\|w\|_{L^{p}(\Omega)}$
$\qquad+\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}\big).$
Here $\tilde{\Lambda}\subset\Lambda$ is introduced in the assumption (6).
Proof.
According to the assumption (6), it is easy to see that for any $\boldsymbol{x}\in\Omega\setminus E_{0}$ and for any $\alpha\in\Lambda$, $a_{ij}^{\alpha}(\boldsymbol{x}),b_{i}^{\alpha}(\boldsymbol{x}),c^{\alpha}(\boldsymbol{x}),f^{\alpha}(\boldsymbol{x})$ are all finite numbers for any $1\leq i,j\leq d$. Since $B\in[L^{1}(\Omega)]^{d\times d},\boldsymbol{c}\in[L^{1}(\Omega)]^{d},w\in L^{1}(\Omega)$, there is a set $E_{1}\subset\Omega$ with zero $d$-dimensional Lebesgue measure, such that $b_{ij},c_{i},w$ are all finite numbers in $\Omega\setminus E_{1}$ for any $1\leq i,j\leq d$.
We choose $t\in\mathbb{R}$ arbitrarily. It is easy to see
$\{\boldsymbol{x}\in\Omega:\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}](\boldsymbol{x})>t\}$
$=$ $\cup_{\alpha\in\Lambda}\{\boldsymbol{x}\in\Omega:[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}](\boldsymbol{x})>t\}.$
We denote by $S_{t}:=\{\boldsymbol{x}\in\Omega:\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}](\boldsymbol{x})>t\}$.
Therefore, for any $\boldsymbol{x}_{0}\in S_{t}\setminus(E_{0}\cup E_{1})$, there is $\alpha_{0}\in\Lambda$ such that
$[A^{\alpha_{0}}:B+\boldsymbol{b}^{\alpha_{0}}\cdot\boldsymbol{c}+c^{\alpha_{0}}w-f^{\alpha_{0}}](\boldsymbol{x}_{0})>t$
We define $\epsilon_{0}=[A^{\alpha_{0}}:B+\boldsymbol{b}^{\alpha_{0}}\cdot\boldsymbol{c}+c^{\alpha_{0}}w-f^{\alpha_{0}}](\boldsymbol{x}_{0})-t>0$. According to the assumption (6), there is $\tilde{\alpha}_{0}\in\tilde{\Lambda}$ such that
$|(A^{\tilde{\alpha}_{0}}-A^{\alpha_{0}}):B(\boldsymbol{x}_{0})|+|(\boldsymbol{b}^{\tilde{\alpha}_{0}}-\boldsymbol{b}^{\alpha_{0}})\cdot\boldsymbol{c}(\boldsymbol{x}_{0})|$
$\qquad+|(c^{\tilde{\alpha}_{0}}-c^{\alpha_{0}})w(\boldsymbol{x}_{0})|+|(f^{\tilde{\alpha}_{0}}-f^{\alpha_{0}})(\boldsymbol{x}_{0})|<\frac{1}{2}\epsilon_{0}.$
So we have
$[A^{\tilde{\alpha}_{0}}:B+\boldsymbol{b}^{\tilde{\alpha}_{0}}\cdot\boldsymbol{c}+c^{\tilde{\alpha}_{0}}w-f^{\tilde{\alpha}_{0}}](\boldsymbol{x}_{0})-t>\epsilon_{0}-\frac{1}{2}\epsilon_{0}>0.$
Therefore, we have
$\boldsymbol{x}_{0}\in\{\boldsymbol{x}\in\Omega\setminus(E_{0}\cup E_{1}):\sup_{\tilde{\alpha}\in\tilde{\Lambda}}[A^{\tilde{\alpha}}:B+\boldsymbol{b}^{\tilde{\alpha}}\cdot\boldsymbol{c}+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}](\boldsymbol{x})>t\}.$
Since $\boldsymbol{x}_{0}\in S_{t}\setminus(E_{0}\cup E_{1})$ is chosen arbitrarily, we have
$S_{t}\setminus(E_{0}\cup E_{1})=\{\boldsymbol{x}\in\Omega\setminus(E_{0}\cup E_{1}):\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}](\boldsymbol{x})>t\}$
$=$ $\{\boldsymbol{x}\in\Omega\setminus(E_{0}\cup E_{1}):\sup_{\tilde{\alpha}\in\tilde{\Lambda}}[A^{\tilde{\alpha}}:B+\boldsymbol{b}^{\tilde{\alpha}}\cdot\boldsymbol{c}+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}](\boldsymbol{x})>t\}$
$=$ $\cup_{\tilde{\alpha}\in\tilde{\Lambda}}\{\boldsymbol{x}\in\Omega\setminus(E_{0}\cup E_{1}):[A^{\tilde{\alpha}}:B+\boldsymbol{b}^{\tilde{\alpha}}\cdot\boldsymbol{c}+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}](\boldsymbol{x})>t\}.$
Since $\tilde{\Lambda}$ has at most countably many elements and $E_{0},E_{1}$ have zero $d$-dimensional Lebesgue measure, we have that $S_{t}$ is a measurable subset of $\Omega$. Since $t\in\mathbb{R}$ is chosen arbitrarily, we can conclude
$\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]$
is a measurable function on $\Omega$ and (74) holds.
According to (74), (4,5) and the fact $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$ imply that $\sup_{\alpha\in\Lambda}[A^{\alpha}:B+\boldsymbol{b}^{\alpha}\cdot\boldsymbol{c}+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega)$ and (75) holds. ∎
Lemma 4.2.
We assume that the assumption (6) is satisfied. For any $\alpha\in\Lambda$, let $\varrho^{\alpha}\in L^{\infty}(\Omega)$. We further assume that there are two positive constants $\nu_{0}\leq\nu_{1}$ and $E_{1}\subset\Omega$ with zero $d$-dimensional Lebesgue measure, such that
(76a) $\nu_{0}\leq\varrho^{\alpha}(\boldsymbol{x})\leq\nu_{1},\qquad\forall\boldsymbol{x}\in\Omega\setminus E_{1},\forall\alpha\in\Lambda;$
[…]
$\qquad|\varrho^{\alpha}(\boldsymbol{x})-\varrho^{\tilde{\alpha}}(\boldsymbol{x})|<\epsilon.$
Here $\tilde{\Lambda}\subset\Lambda$ is introduced in the assumption (6).
We choose $w\in W^{2,1}(\Omega)$ arbitrarily. Then $\sup_{\alpha\in\Lambda}\varrho^{\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]$ is a measurable function on $\Omega$. If $\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=0$ in $\Omega$, then
$\sup_{\alpha\in\Lambda}\varrho^{\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=0\qquad\text{ in }\Omega.$
Proof.
According to the assumption (6), it is easy to see that for any $\boldsymbol{x}\in\Omega\setminus E_{0}$ and for any $\alpha\in\Lambda$, $a_{ij}^{\alpha}(\boldsymbol{x}),b_{i}^{\alpha}(\boldsymbol{x}),c^{\alpha}(\boldsymbol{x}),f^{\alpha}(\boldsymbol{x})$ are all finite numbers for any $1\leq i,j\leq d$. Then the assumption (6) is satisfied by $\left(\varrho^{\alpha}A^{\alpha},\varrho^{\alpha}\boldsymbol{b}^{\alpha},\varrho^{\alpha}c^{\alpha},\varrho^{\alpha}f^{\alpha}\right)$ with $\tilde{\Lambda}\subset\Lambda$ and the set $E_{0}\cup E_{1}$. So Lemma 4.1 implies that $\sup_{\alpha\in\Lambda}\varrho^{\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}u-f^{\alpha}]$ is a measurable function on $\Omega$.
Since $w\in W^{2,1}(\Omega)$, there is a set $E_{1}^{\prime}\subset\Omega$ with zero $d$-dimensional Lebesgue measure such that $\frac{\partial^{2}w}{\partial x_{i}\partial x_{j}}(\boldsymbol{x}),\frac{\partial w}{\partial x_{i}}(\boldsymbol{x}),w(\boldsymbol{x})$ are all finite numbers for any $\boldsymbol{x}\in\Omega\setminus E_{1}^{\prime}$ and for any $1\leq i,j\leq d$. We choose $\boldsymbol{x}_{0}\in\Omega\setminus(E_{0}\cup E_{1}\cup E_{1}^{\prime})$ arbitrarily. We claim that if $\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}](\boldsymbol{x}_{0})=0$, then
$\sup_{\alpha\in\Lambda}\varrho^{\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}](\boldsymbol{x}_{0})=0.$
The proof will be complete if this claim is true.
We prove this claim in the following. The fact $\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}](\boldsymbol{x}_{0})=0$ is equivalent to
(77a) $[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}](\boldsymbol{x}_{0})\leq 0,\qquad\forall\alpha\in\Lambda;$
(77b) $\text{there is a sequence }\{\alpha_{i}\}_{i=1}^{+\infty}\subset\Lambda\text{ such that }$
$\qquad\lim_{i\rightarrow+\infty}[A^{\alpha_{i}}:D^{2}w+\boldsymbol{b}^{\alpha_{i}}\cdot\nabla w+c^{\alpha_{i}}w-f^{\alpha_{i}}](\boldsymbol{x}_{0})=0.$
By (76a,77a), we have
$\varrho^{\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}](\boldsymbol{x}_{0})\leq 0,\qquad\forall\alpha\in\Lambda.$
On the other hand, (76a,77a, 77b) implies
$\lim_{i\rightarrow+\infty}\varrho^{\alpha}[A^{\alpha_{i}}:D^{2}w+\boldsymbol{b}^{\alpha_{i}}\cdot\nabla w+c^{\alpha_{i}}w-f^{\alpha_{i}}](\boldsymbol{x}_{0})=0.$
So the claim has been proven. Therefore, the proof is complete. ∎
The following Lemma 4.3 shows that the assumptions on coefficients and source term of (3) used in [40] and ([23, Proposition $4.3$] imply the assumption (6).
Lemma 4.3.
We assume that for any $\epsilon>0$, there is $\delta>0$ such that
(78) $|A^{\alpha}(\boldsymbol{x})-A^{\alpha}(\boldsymbol{y})|+|\boldsymbol{b}^{\alpha}(\boldsymbol{x})-\boldsymbol{b}^{\alpha}(\boldsymbol{y})|+|c^{\alpha}(\boldsymbol{x})-c^{\alpha}(\boldsymbol{y})|+|f^{\alpha}(\boldsymbol{x})-f^{\alpha}(\boldsymbol{y})|<\epsilon,$
for any $\alpha\in\Lambda$ and for any $\boldsymbol{x},\boldsymbol{y}\in\Omega$ satisfying $|\boldsymbol{x}-\boldsymbol{y}|<\delta$. Then the assumption (6) is satisfied.
Proof.
Without losing of generality, we assume $A^{\alpha},\boldsymbol{b}$ and $c^{\alpha}$ are zero matrices, zero vectors and zero for any $\alpha\in\Lambda$, respectively.
Let $n$ be an arbitrary positive integer. We define $\Omega_{n}=\Omega\cap\{(l_{1}2^{-n},\cdots,l_{d}2^{-n}):(l_{1},\cdots,l_{d})\in\mathbb{Z}^{d}\}$. We define a countable subset $\Lambda_{n}$ of $\Lambda$. For any $\boldsymbol{y}\in\Omega_{n}$ and any $l\ in\mathbb{Z}$, if there is $\alpha\in\Lambda$ such that $l2^{-n}\leq f^{\alpha}<(l+1)2^{-n}$, then we put this $\alpha$ into $\Lambda_{n}$. Notice that though there may be more than one $\alpha\in\Lambda$ qualified, we only choose at most one $\alpha$ in the above procedure. Therefore, $\Lambda_{n}$ has at most countably many elements. We define $\tilde{\Lambda}=\cup_{n=1}^{\infty}\Lambda_{n}\subset\tilde{\Lambda}$.
We choose $\boldsymbol{x}\in\Omega$ and $\alpha\in\Lambda$ and $\epsilon>0$ arbitrarily. Then there is a sequence $\{\boldsymbol{y}_{n}\}_{n=1}^{\infty}\subset\tilde{\Lambda}$ such that $\lim_{n\rightarrow+\infty}|\boldsymbol{y}_{n}-\boldsymbol{x}|=0$ and $\boldsymbol{y}_{n}\in\Lambda_{n}$ for any positive integer $n$. Then by (78), there is a positive integer $N_{0}$ such that
$|f^{\beta}(\boldsymbol{y}_{n})-f^{\beta}(\boldsymbol{x})|<\frac{\epsilon}{3},\qquad 2^{-n}<\frac{\epsilon}{3},$
for any $n\geq N_{0}$ and any $\beta\in\Lambda$.
There is an integer $l_{0}$ such that $l2^{-N_{0}}\leq f^{\alpha}(\boldsymbol{y}_{N_{0}})<(l+1)2^{-N_{0}}$. By the construction of $\Lambda_{n}\subset\tilde{\alpha}$, there is $\tilde{\alpha}\in\Lambda_{n}$ such that
$l_{0}2^{-N_{0}}\leq f^{\tilde{\alpha}}(\boldsymbol{y}_{N_{0}})<(l_{0}+1)2^{-N_{0}}.$
Then we have
$|f^{\tilde{\alpha}}(\boldsymbol{y}_{N_{0}})-f^{\alpha}(\boldsymbol{y}_{N_{0}})|<2^{-N_{0}}<\frac{\epsilon}{3}.$
We notice that
$|f^{\alpha}(\boldsymbol{y}_{n})-f^{\alpha}(\boldsymbol{x})|<\frac{\epsilon}{3},\qquad|f^{\tilde{\alpha}}(\boldsymbol{y}_{n})-f^{\tilde{\alpha}}(\boldsymbol{x})|<\frac{\epsilon}{3}.$
Then we have that
$|f^{\alpha}(\boldsymbol{x})-f^{\tilde{\alpha}}(\boldsymbol{x})|<\epsilon.$
We can conclude that the proof is complete. ∎
4.1. Well-posedness of Hamilton-Jacobi-Bellman equation
Theorem 4.1.
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (4,5) and the assumption (6) hold. There is a constant $0\leq\underline{\kappa}<1$ which may depend on $p$, such that if
(79) $\dfrac{|A^{\alpha}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\alpha}(\boldsymbol{x})|^{2}+\lambda^{-2}|c^{\alpha}(\boldsymbol{x})|^{2}}{\left(\text{Tr}A^{\alpha}(\boldsymbol{x})+\lambda^{-1}|c^{\alpha}(\boldsymbol{x})|\right)^{2}}$
$\leq$ $\dfrac{1}{d+\epsilon},\qquad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},\forall\alpha\in\Lambda,$
for some $\epsilon\in[\underline{\kappa},1)$ and $\lambda>0$, then there is a unique solution $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfying (3). Furthermore, there is a positive constant $C$ which may depend on $p$, such that
$\|u\|_{W^{2,p}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
Here $\tilde{\Lambda}\subset\Lambda$ is introduced in the assumption (6). $p$ is valid in the range described in (1a). In addition, $\underline{\kappa}=0$ if $p=2$ and the domain is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$).
Remark 4.2.
In special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$, Theorem 4.1 still holds if the condition (79) is replaced by
(80) $\dfrac{|A^{\alpha}(\boldsymbol{x})|^{2}}{\left(\text{Tr}A^{\alpha}(\boldsymbol{x})\right)^{2}}\leq\dfrac{1}{d-1+\epsilon},\qquad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},\forall\alpha\in\Lambda,$
for some $\epsilon\in[\underline{\kappa}^{\prime},1)$ where $\underline{\kappa}^{\prime}\in[0,1)$. In addition, $\underline{\kappa}^{\prime}=0$ if $p=2$ and the domain is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). The proof is similar. We would like to point out that the condition (79) and the condition (80) are exactly the Cordès conditions in [40].
Proof.
Let $p$ be any number in the range described in this Theorem.
We denote by
$\|v\|_{W_{\lambda}^{2,p}(\Omega)}=\big(\|D^{2}v\|_{L^{p}(\Omega)}^{p}+2\lambda\|\nabla v\|_{L^{p}(\Omega)}^{p}+\lambda^{2}\|v\|_{L^{p}(\Omega)}^{p}\big)^{1/p},\qquad\forall v\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega).$
According to the assumption (6), it is easy to see that for any $\boldsymbol{x}\in\Omega\setminus E_{0}$ and for any $\alpha\in\Lambda$, $a_{ij}^{\alpha}(\boldsymbol{x}),b_{i}^{\alpha}(\boldsymbol{x}),c^{\alpha}(\boldsymbol{x}),f^{\alpha}(\boldsymbol{x})$ are all finite numbers for any $1\leq i,j\leq d$.
For any $\alpha\in\Lambda$, we define a function $\gamma^{\lambda,\alpha}$ on $\Omega$ such that for any $\boldsymbol{x}\in\Omega$ almost everywhere,
(81) $\gamma^{\lambda,\alpha}(\boldsymbol{x})=\dfrac{\Sigma_{1\leq i\leq d}a_{ii}^{\alpha}(\boldsymbol{x})+\lambda^{-1}|c^{\alpha}(\boldsymbol{x})|}{|A^{\alpha}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\alpha}(\boldsymbol{x})|^{2}+(\lambda^{-1}c^{\alpha}(\boldsymbol{x}))^{2}}.$
Here $|B|=\big(\Sigma_{1\leq i,j\leq d}|b_{ij}|^{2}\big)^{1/2}$ for any $B\in\mathbb{R}^{d\times d}$ and $|\boldsymbol{\beta}|=\big(\Sigma_{1\leq i\leq d}|\beta_{i}|^{2}\big)^{1/2}$ for any $\boldsymbol{\beta}\in\mathbb{R}^{d}$.
By (4,5) and the fact $\tilde{\Lambda}$ is a countable set, there is a set $E_{1}\supset E_{0}$ with zero $d$-dimensional Lebesgue measure such that
$\gamma^{\lambda,\tilde{\alpha}}(\boldsymbol{x})=\dfrac{\Sigma_{1\leq i\leq d}a_{ii}^{\tilde{\alpha}}(\boldsymbol{x})+\lambda^{-1}|c^{\tilde{\alpha}}(\boldsymbol{x})|}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|^{2}+(\lambda^{-1}c^{\tilde{\alpha}}(\boldsymbol{x}))^{2}},$
$C_{1}\leq\gamma^{\lambda,\tilde{\alpha}}(\boldsymbol{x})\leq C_{2},$
for any $\boldsymbol{x}\in\Omega\setminus E_{1}$ and any $\tilde{\alpha}\in\tilde{\Lambda}$. Here $0<C_{1}\leq C_{2}<+\infty$ are two constants independent of $\boldsymbol{x}\in\Omega\setminus E_{1}$ and $\tilde{\alpha}\in\tilde{\Lambda}$. Then by the assumption (6), we have
$\gamma^{\lambda,\alpha}(\boldsymbol{x})=\dfrac{\Sigma_{1\leq i\leq d}a_{ii}^{\alpha}(\boldsymbol{x})+\lambda^{-1}|c^{\alpha}(\boldsymbol{x})|}{|A^{\alpha}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\alpha}(\boldsymbol{x})|^{2}+(\lambda^{-1}c^{\alpha}(\boldsymbol{x}))^{2}},$
$C_{1}\leq\gamma^{\lambda,\alpha}(\boldsymbol{x})\leq C_{2},$
for any $\boldsymbol{x}\in\Omega\setminus E_{1}$ and any $\alpha\in\Lambda$.
Obviously, the assumption (6) is satisfied by $(\gamma^{\lambda,\alpha}A^{\alpha},\gamma^{\lambda,\alpha}\boldsymbol{b}^{\alpha},\gamma^{\lambda,\alpha}c^{\alpha},\gamma^{\lambda,\alpha}f^{\alpha})$ with $\tilde{\Lambda}\subset\Lambda$ and the set $E_{1}$. Then by Lemma 4.1 and (4,5), we have that for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
(82) $\sup_{\alpha\in\Lambda}\gamma^{\lambda,\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]\in L^{p}(\Omega),$
$\|\sup_{\alpha\in\Lambda}\gamma^{\lambda,\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]\|_{L^{p}(\Omega)}$
$\leq$ $C\big(\|w\|_{W^{2,p}(\Omega)}+\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}\big).$
Then Lemma 4.2 implies that for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$, $\sup_{\alpha\in\Lambda}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=0$ in $\Omega$ is equivalent to
$\sup_{\alpha\in\Lambda}\gamma^{\lambda,\alpha}[A^{\alpha}:D^{2}w+\boldsymbol{b}^{\alpha}\cdot\nabla w+c^{\alpha}w-f^{\alpha}]=0\quad\text{ in }\Omega,$
which is equivalent to
(83) $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}w+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}]=0\quad\text{ in }\Omega,$
due to (74).
We define a mapping $M_{\lambda}:L^{p}(\Omega)\rightarrow W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ by
$(\Delta-\lambda)(M_{\lambda}g)=g\text{ in }\Omega,$
for any $g\in L^{p}(\Omega)$. According to Theorem 3.4, we have that for any $g\in L^{p}(\Omega)$,
(84) $\|M_{\lambda}g\|_{W_{\lambda}^{2,p}(\Omega)}\leq\underline{C}_{p,\lambda}\|g\|_{L^{p}(\Omega)},$
where the constant $\underline{C}_{p,\lambda}$ may depend on $p$ and $\lambda$. In fact, $\underline{C}_{2,\lambda}=1$ for any $\lambda\geq 1$ with convex domain due to [40, Theorem $2$].
By (83) and the definition of $M_{\lambda}$, $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfies the Hamilton-Jacobi-Bellman equation (3) if and only if
$-u=M_{\lambda}\big(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}u+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u+c^{\tilde{\alpha}}u-f^{\tilde{\alpha}}]-(\Delta u-\lambda u)\big).$
On the other hand, (82) and (84) imply that
(85) $w\rightarrow T_{\lambda}(w):=M_{\lambda}\big(\sup_{\tilde{\alpha}\in\Lambda}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}w+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w+c^{\tilde{\alpha}}w-f^{\tilde{\alpha}}]-(\Delta-\lambda)w\big)$
is a well defined mapping from $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ into itself. Therefore, in order to to the existence and uniqueness of the solution of (3), it is sufficient to prove that the mapping (85) is a contraction with respect to the norm $\|\cdot\|_{W_{\lambda}^{2,p}(\Omega)}$.
By (79) and the proof of [40, Lemma $1$] ($\underline{\kappa}$ will be determined later), we have that for any $w_{1},w_{2}\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
(86) $\|\sup_{\tilde{\alpha}\in\Lambda}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}w_{1}+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{1}+c^{\tilde{\alpha}}w_{1}-f^{\tilde{\alpha}}$]
$\qquad-\sup_{\tilde{\alpha}\in\Lambda}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}w_{2}+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{2}+c^{\tilde{\alpha}}w_{2}-f^{\tilde{\alpha}}]-(\Delta-\lambda)(w_{1}-w_{2})\|_{L^{p}(\Omega)}$
$\leq$ $\underline{C}_{p,\lambda}^{\prime}\sqrt{1-\epsilon}\|(w_{1}-w_{2})\|_{W_{\lambda}^{2,p}(\Omega)}.$
In fact, $\underline{C}_{2,\lambda}^{\prime}=1$ with convex domain.
By (84,86), we have
[…]
$\qquad-M_{\lambda}\big(\sup_{\tilde{\alpha}\in\Lambda}\gamma^{\lambda,\tilde{\alpha}}[A^{\tilde{\alpha}}:D^{2}w_{2}+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{2}+c^{\tilde{\alpha}}w_{2}-f^{\tilde{\alpha}}]\big)-(w_{1}-w_{2})\|_{L^{p}(\Omega)}$
$\leq$ $\underline{C}_{p,\lambda}\underline{C}_{p,\lambda}^{\prime}\sqrt{1-\epsilon}\|(w_{1}-w_{2})\|_{W_{\lambda}^{2,p}(\Omega)}.$
We can choose
(88) $\underline{\kappa}=1-\frac{1}{(\underline{C}_{2,\lambda}\underline{C}_{2,\lambda}^{\prime})^{2}}.$
By (84) and (86), if (79) is satisfied with $\epsilon\in(\underline{\kappa},1)$, $T_{\lambda}$ is a contraction. Then there is a unique solution $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfying the Hamilton-Jacobi-Bellman equation (3) which is equivalent to (83). Since $\underline{C}_{2,\lambda}=\underline{C}_{2,\lambda}^{\prime}=1$, $\underline{\kappa}=1$ if $p=2$ with convex domain.
By (87) with $w_{1}=u$ and $w_{2}=0$, we have
$\|u-M_{\lambda}(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}f^{\tilde{\alpha}})\|_{W_{\lambda}^{2,p}(\Omega)}\leq\underline{C}_{p,\lambda}\underline{C}_{p,\lambda}^{\prime}\sqrt{1-\epsilon}\|u\|_{W_{\lambda}^{2,p}(\Omega)}.$
Since $\underline{C}_{p,\lambda}\underline{C}_{p,\lambda}^{\prime}\sqrt{1-\epsilon}<1$, we have
$\|u\|_{W^{2,p}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
Therefore, the proof is complete. ∎
4.2. Optimal convergence of numerical solution of FEM for Hamilton-Jacobi-Bellman equation
Theorem 4.3.
Let $\Omega$ be a Lipschitz polyhedral in $\mathbb{R}^{d}$ ($d=2,3$). We assume that $\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\in L^{p}(\Omega)$, and (4,5) and the assumption (6) hold. There is a constant $0\leq\kappa<1$ which my depend on $p$, such that if
(89) $\dfrac{|A^{\alpha}(\boldsymbol{x})|^{2}(\text{Tr}A^{\alpha}(\boldsymbol{x}))^{-2}}{1+2\lambda^{-1}|c^{\alpha}(\boldsymbol{x})|(\text{Tr}A^{\alpha}(\boldsymbol{x}))^{-1}-|A^{\alpha}(\boldsymbol{x})|^{-2}\big(\lambda^{-2}|c^{\alpha}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\alpha}(\boldsymbol{x})|^{2}\big)}$
$\leq$ $\dfrac{1}{d+\epsilon},\quad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},\forall\alpha\in\Lambda,$
for some constants $\epsilon\in[\kappa,1)$ and $\lambda>0$, then there is a unique numerical solution $u_{h}\in V_{h}$ of the finite element method (16) for any $0<h<h_{3}$. In addition, there is a positive constant $C$ which may depend on $p$, such that
$\|u_{h}\|_{W^{2,p}_{h}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
Here $\tilde{\Lambda}\subset\Lambda$ is introduced in the assumption (6). Furthermore, if there are $\epsilon\in[0,1)$ and $\lambda>0$ such that (89) holds, then (79) in Theorem 4.1 holds as well for the same $\epsilon$ and $\lambda$. We also have the convergent result that for any $\chi_{h}\in V_{h}$,
(90) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
if $0<h<h_{3}$. Here $\overline{P}_{h}$ is the standard $L^{2}$-orthogonal projection onto $\overline{V}_{h}]^{d\times d}$. $p$ is valid in the range described in (1b).
Remark 4.4.
In special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$, Theorem 4.3 still holds (no $\lambda$ any more) if the condition (79) is replaced by
(91) $\dfrac{|A^{\alpha}(\boldsymbol{x})|^{2}}{\left(\text{Tr}A^{\alpha}(\boldsymbol{x})\right)^{2}}\leq\dfrac{1}{d-1+\epsilon},\qquad\forall\boldsymbol{x}\in\Omega\text{ almost everywhere},\forall\alpha\in\Lambda,$
for some $\epsilon\in[\kappa^{\prime},1)$ where $\kappa^{\prime}\in[0,1)$. The proof is much simpler. We would like to emphasize that even for $p=2$, $\kappa^{\prime}$ is not known explicitly and probably bigger than 0, since the constant of the discrete Miranda–Talenti inequality of the discrete Lapalcian operator $L_{I,h}$ associated with the FEM (16) is not known explicitly and probably bigger than $1$.
Remark 4.5.
(The condition (89) includes non-trivial example) We investigate that under which restrictions, the condition (89) can be valid for [40, Example $1$].
We recall that [40, Example $1$] is defined in two dimensional domain, such that
$|A^{\alpha}|^{2}=\dfrac{1+(\sin\theta)^{2}}{2},\quad\text{Tr}A^{\alpha}=1,\quad\boldsymbol{b}^{\alpha}=\boldsymbol{0},\quad c^{\alpha}=-c_{0}\text{ with }c_{0}>0,$
for any $\alpha\in\Lambda$ with $\theta\in[0,\frac{\pi}{3}]$. Then the condition (89) is equivalent to
(92) $\dfrac{\frac{1}{2}(1+(\sin\theta)^{2})}{1+2(\frac{c_{0}}{\lambda})-\frac{2}{1+(\sin\theta)^{2}}(\frac{c_{0}}{\lambda})^{2}}\leq\frac{1}{2+\epsilon}.$
It is easy to verify that (92) holds for any $\epsilon\in(0,1)$ if $\theta=0$ and $\lambda=2c_{0}$, and holds also for any $\epsilon\in(0,\frac{1}{7}]$ if $\theta=\frac{\pi}{3}$ and $\lambda=\frac{8}{7}c_{0}$.
We notice that $\epsilon\in(\kappa,1)$ where $\kappa\in[0,1)$ is not known explicitly. Then by the continuity argument, there exists $\beta\in(0,\frac{\pi}{3}]$ such that (92) with $\epsilon=\kappa$ holds for some $\lambda\in(\frac{8}{7}c_{0},2c_{0})$ and for any $\theta\in[0,\beta]$. Therefore, we can conclude that conservatively speaking, the condition (89) can be satisfied by [40, Example $1$] with probably more restricted region of the parameter $\theta$ (a non-trivial closed subinterval contained in $0,\frac{\pi}{3}]$).
Proof.
Let $p$ be any number in the range described in this Theorem.
Since (4,5) and the assumption (6) hold, it is obvious to mimic the proof of Theorem 4.1 (the argument to obtain (83)) to have that the finite element method (16) is equivalent to find $u_{h}\in V_{h}$ satisfying
(93) $(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla u_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u_{h}+c^{\tilde{\alpha}}u_{h}-f^{\tilde{\alpha}}],v_{h})_{\Omega}=0,\qquad\forall v_{h}\in V_{h}.$
We denote by
$\|v_{h}\|_{W_{h,\lambda}^{2,p}(\Omega)}=\big(\|v_{h}\|_{W_{h}^{2,p}(\Omega)}^{p}+2\lambda\|\nabla v_{h}\|_{L^{p}(\Omega)}^{p}+\lambda^{2}\|v_{h}\|_{L^{p}(\Omega)}^{p}\big)^{1/p},\qquad\forall v_{h}\in V_{h}.$
For any $w_{h}\in V_{h}$, we define $F_{h}(w_{h})\in V_{h}$ by
$(F_{h}(w_{h}),v_{h})_{\Omega}$
$=$ $(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla u_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u_{h}+c^{\tilde{\alpha}}u_{h}-f^{\tilde{\alpha}}],v_{h})_{\Omega},\quad\forall v_{h}\in V_{h}.$
We define a mapping $M_{\lambda,h}:V_{h}\rightarrow V_{h}$ such that for any $g_{h}\in V_{h}$,
$-(\nabla(M_{\lambda,h}(g_{h})),\nabla v_{h})_{\Omega}-\lambda(M_{\lambda,h}(g_{h}),v_{h})_{\Omega}=(g_{h},v_{h})_{\Omega},\qquad\forall v_{h}\in V_{h}.$
It is easy to see that for any $w_{h}\in V_{h}$,
$M_{\lambda,h}(\mathcal{L}_{I_{d},h}w_{h}-\lambda w_{h})=w_{h},$
where $\mathcal{L}_{I_{d},h}w_{h}\in V_{h}$ satisfying $(\mathcal{L}_{I_{d},h}w_{h},v_{h})_{\Omega}=(I_{d}:\overline{\nabla}_{h}(\nabla w_{h}),v_{h})_{\Omega}$ for any $v_{h}\in V_{h}$. Then by Theorem 3.3, we have that
(94) $\|M_{\lambda,h}(g_{h})\|_{W_{h,\lambda}^{2,p}(\Omega)}\leq C_{p,\lambda}\|g_{h}\|_{L^{p}(\Omega)},\qquad\forall g_{h}\in V_{h},\forall 0<h<h_{3},$
where the constant $C_{p,\lambda}$ may depend on $p$ and $\lambda$.
We notice that $u_{h}\in V_{h}$ is a numerical solution of the finite element method (93) if and only if $F_{h}(u_{h})=0$ in $\Omega$, which is equivalent to
$-u_{h}=M_{\lambda,h}\big(F_{h}(u_{h})-(\mathcal{L}_{I_{d},h}u_{h}-\lambda u_{h})\big).$
Therefore, to prove the existence and uniqueness of (93), it is sufficient to prove the mapping
(95) $w_{h}\rightarrow M_{\lambda,h}\big(F_{h}(w_{h})-(\mathcal{L}_{I_{d},h}w_{h}-\lambda w_{h})\big)$
is a contraction from $V_{h}$ into itself with respect to the $\|\cdot\|_{W_{h,\lambda}^{2,p}(\Omega)}$ norm.
We choose $w_{1,h},w_{2,h}\in V_{h}$ arbitrarily. We notice that for any $v_{h}\in V_{h}$
$F_{h}(v_{h})-(\mathcal{L}_{I_{d},h}v_{h}-\lambda v_{h})$
$=$ $P_{h}\big(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla v_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla v_{h}+c^{\tilde{\alpha}}v_{h}-f^{\tilde{\alpha}}]-(\overline{\nabla}_{h}(\nabla v_{h})-\lambda v_{h})\big).$
Here $P_{h}$ is the standard $L^{2}$-orthogonal projection onto $V_{h}$. Therefore, we have
(96) $\|\big(F_{h}(w_{1,h})-(\mathcal{L}_{I_{d},h}w_{1,h}-\lambda w_{1,h})\big)$
$\qquad-\big(F_{h}(w_{2,h})-(\mathcal{L}_{I_{d},h}w_{2,h}-\lambda w_{2,h})\big)\|_{L^{p}(\Omega)}$
$\leq$ $C_{p}^{\prime}\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla w_{1,h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{1,h}+c^{\tilde{\alpha}}w_{1,h}-f^{\tilde{\alpha}}$]
$\qquad-\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla w_{2,h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{2,h}+c^{\tilde{\alpha}}w_{2,h}-f^{\tilde{\alpha}}$]
$\qquad-(\overline{\nabla}_{h}(\nabla(w_{1,h}-w_{2,h}))-\lambda(w_{1,h}-w_{2,h}))\|_{L^{p}(\Omega)}$
According to the proof of [40, Lemma $1$], we have that for any $\boldsymbol{x}\in\Omega$ almost everywhere,
(97) $|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla w_{1,h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{1,h}+c^{\tilde{\alpha}}w_{1,h}-f^{\tilde{\alpha}}](\boldsymbol{x})$
$\qquad-\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla w_{2,h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla w_{2,h}+c^{\tilde{\alpha}}w_{2,h}-f^{\tilde{\alpha}}](\boldsymbol{x})$
$\qquad-(\overline{\nabla}_{h}(\nabla(w_{1,h}-w_{2,h}))-\lambda(w_{1,h}-w_{2,h}))(\boldsymbol{x})|$
$\leq$ $\big(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\sqrt{E^{\tilde{\alpha}}(\boldsymbol{x})}\big)\sqrt{|\overline{\nabla}_{h}(\nabla w_{h}(\boldsymbol{x}))|^{2}+2\lambda|\nabla w_{h}(\boldsymbol{x})|^{2}+\lambda^{2}|w_{h}(\boldsymbol{x})|^{2}},$
where $w_{h}=w_{1,h}-w_{2,h}$ and
$E^{\tilde{\alpha}}(\boldsymbol{x})=$ $|\gamma^{\tilde{\alpha}}(\boldsymbol{x})A^{\tilde{\alpha}}(\boldsymbol{x})-I_{d}|^{2}+|\gamma^{\tilde{\alpha}}(\boldsymbol{x})|^{2}\frac{|\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{2\lambda}+\dfrac{|\lambda+c^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{\lambda^{2}}$
$=$ $\frac{|\gamma^{\tilde{\alpha}}(\boldsymbol{x})|^{2}|c^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{\lambda^{2}}+\gamma^{\tilde{\alpha}}(\boldsymbol{x})\big(\frac{\gamma^{\tilde{\alpha}}(\boldsymbol{x})|\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|}{2\lambda}-\frac{2|c^{\tilde{\alpha}}(\boldsymbol{x})|}{\lambda}\big)$
$\qquad+\big(d+1+|\gamma^{\tilde{\alpha}}(\boldsymbol{x})|^{2}|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}-2\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x})\big).$
The last equation above is due to the fact $c^{\tilde{\alpha}}(\boldsymbol{x})\leq 0$ for any $\tilde{\alpha}\in\tilde{\Lambda}$. By the definition of $\gamma^{\tilde{\alpha}}$ (see (15)),
$E^{\tilde{\alpha}}(\boldsymbol{x})=$ $\frac{(\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x}))^{2}|c^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{4}}\lambda^{-2}+\frac{\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x})}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}\big(\frac{\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x})}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}\frac{|\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{2}-2|c^{\tilde{\alpha}}(\boldsymbol{x})|\big)\lambda^{-1}$
$\qquad+(d+1-\frac{(\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x}))^{2}}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}).$
Then for any $\epsilon\in(0,1]$, $E^{\tilde{\alpha}}(\boldsymbol{x})\leq 1-\epsilon$ is equivalent to
$-\frac{(\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x}))^{2}|c^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{4}}\lambda^{-2}-\frac{\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x})}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}\big(\frac{\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x})}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}\frac{|\boldsymbol{b}^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}{2}-2|c^{\tilde{\alpha}}(\boldsymbol{x})|\big)\lambda^{-1}$
$\qquad+\frac{(\text{Tr}A^{\tilde{\alpha}}(\boldsymbol{x}))^{2}}{|A^{\tilde{\alpha}}(\boldsymbol{x})|^{2}}\geq d+\epsilon,$
which is equivalent to
$\dfrac{|A^{\alpha}(\boldsymbol{x})|^{2}(\text{Tr}A^{\alpha}(\boldsymbol{x}))^{-2}}{1+2\lambda^{-1}|c^{\alpha}(\boldsymbol{x})|(\text{Tr}A^{\alpha}(\boldsymbol{x}))^{-1}-|A^{\alpha}(\boldsymbol{x})|^{-2}\big(\lambda^{-2}|c^{\alpha}(\boldsymbol{x})|^{2}+(2\lambda)^{-1}|\boldsymbol{b}^{\alpha}(\boldsymbol{x})|^{2}\big)}\leq\dfrac{1}{d+\epsilon}.$
By (89), we have that for any $\epsilon\in(\kappa,1)$,
(98) $E^{\tilde{\alpha}}(\boldsymbol{x})\leq$ $1-\epsilon,$
for any $\boldsymbol{x}\in\Omega$ almost everywhere and for any $\tilde{\alpha}\in\tilde{\Lambda}$. Here the constant $\kappa\in[0,1)$ will be determined later.
By (96,97,98), we have
(99) $\|\big(F_{h}(w_{1,h})-(\mathcal{L}_{I_{d},h}w_{1,h}-\lambda w_{1,h})\big)$
$\qquad-\big(F_{h}(w_{2,h})-(\mathcal{L}_{I_{d},h}w_{2,h}-\lambda w_{2,h})\big)\|_{L^{p}(\Omega)}$
$\leq$ $C_{p,\lambda}^{\prime}C_{p}^{\prime}\sqrt{1-\epsilon}\big(\|\overline{\nabla}_{h}(\nabla(w_{1,h}-w_{2,h}))\|_{L^{p}(\Omega)}^{p}$
$\qquad+2\lambda\|\nabla(w_{1,h}-w_{2,h})\|_{L^{p}(\Omega)}^{p}+\lambda^{2}\|w_{1,h}-w_{2,h}\|_{L^{p}(\Omega)}^{p}\big)^{1/p}.$
By (2.1), it is easy to see that
(100) $\|\overline{\nabla}_{h}(\nabla(w_{1,h}-w_{2,h}))\|_{L^{p}(\Omega)}^{p}\leq C_{p}^{\prime\prime}\|w_{1,h}-w_{2,h}\|_{W_{h}^{2,p}(\Omega)}.$
By (94, 99, 100), we have
(101) $\|M_{\lambda,h}\big(F_{h}(w_{1,h})-(\mathcal{L}_{I_{d},h}w_{1,h}-\lambda w_{1,h})\big)$
$\qquad-M_{\lambda,h}\big(F_{h}(w_{2,h})-(\mathcal{L}_{I_{d},h}w_{2,h}-\lambda w_{2,h})\big)\|_{W_{h,\lambda}^{2.p}(\Omega)}$
$\leq$ $\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}\|w_{1,h}-w_{2,h}\|_{W_{h,\lambda}^{2,p}(\Omega)}.$
In order to make the mapping (95) a contraction, it is sufficient to choose
(102) $\kappa=1-\dfrac{1}{\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)^{2}},$
such that $\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}<1$ for any $\epsilon\in(\kappa,1]$. Therefore, the existence and uniqueness of numerical solution of the finite element method (16) is proven.
Let $u_{h}\in V_{h}$ be the numerical solution of (16). Then we have
$F_{h}(u_{h})=0\text{ and }u_{h}=-M_{\lambda,h}\big(F_{h}(u_{h})-(\mathcal{L}_{I_{d},h}u_{h}-\lambda u_{h})\big)\text{ in }\Omega.$
By (101) with $w_{1,h}=u_{h}$ and $w_{2,h}=0$, we have
$\|u_{h}-M_{\lambda,h}\big(\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}f^{\tilde{\alpha}}\big)\|_{W_{h,\lambda}^{2,p}(\Omega)}\leq\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}\|u_{h}\|_{W_{h,\lambda}^{2,p}(\Omega)}.$
Since $\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}<1$ for any $\epsilon\in(\kappa,1]$, we have that
(103) $\|u_{h}\|_{W^{2,p}_{h}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
It is straightforward to verify that if there are $\epsilon\in(0,1]$ and $\lambda>0$ such that (89) holds, then (79) in Theorem 4.1 holds as well for the same $\epsilon$ and $\lambda$. Furthermore, it is easy to check that $\kappa\geq\underline{\kappa}$ which is introduced in (88). Then by Theorem 4.1, there is a unique solution $u\in W_{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ of the Hamilton-Jacobi-Bellman equation (3).
We choose $\chi_{h}\in V_{h}$ arbitrarily. By (101) with $w_{1,h}=u_{h}$ and $w_{2,h}=\chi_{h}$, we have
$\|M_{\lambda,h}(F_{h}(\chi_{h}))+(u_{h}-\chi_{h})\|_{W_{h,\lambda}^{2,p}(\Omega)}\leq\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}\|u_{h}-\chi_{h}\|_{W_{h,\lambda}^{2,p}(\Omega)}.$
By (94) and the fact $\big(C_{p,\lambda}C_{p,\lambda}^{\prime}C_{p}^{\prime}C_{p}^{\prime\prime}\big)\sqrt{1-\epsilon}<1$, we have
$\|u_{h}-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|F_{h}(\chi_{h})\|_{L^{p}(\Omega)}.$
By the same argument in the proof of Theorem 4.1, it is easy to see that
$\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\alpha}[A^{\tilde{\alpha}}:D^{2}u+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u+c^{\tilde{\alpha}}u-f^{\tilde{\alpha}}]=0\text{ in }\Omega.$
We notice that
$\|F_{h}(\chi_{h})\|_{L^{p}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla\chi_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla\chi_{h}+c^{\tilde{\alpha}}\chi_{h}-f^{\tilde{\alpha}}]\|_{L^{p}(\Omega)}$
$=$ $C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla\chi_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla\chi_{h}+c^{\tilde{\alpha}}\chi_{h}-f^{\tilde{\alpha}}$]
$\qquad-\sup_{\tilde{\alpha}\in\tilde{\Lambda}}\gamma^{\alpha}[A^{\tilde{\alpha}}:D^{2}u+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u+c^{\tilde{\alpha}}u-f^{\tilde{\alpha}}]\|_{L^{p}(\Omega)}$
$\leq$ $C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|\gamma^{\tilde{\alpha}}[A^{\tilde{\alpha}}:\overline{\nabla}_{h}(\nabla\chi_{h})+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla\chi_{h}+c^{\tilde{\alpha}}\chi_{h}-f^{\tilde{\alpha}}$]
$\qquad-\gamma^{\alpha}[A^{\tilde{\alpha}}:D^{2}u+\boldsymbol{b}^{\tilde{\alpha}}\cdot\nabla u+c^{\tilde{\alpha}}u-f^{\tilde{\alpha}}]|\|_{L^{p}(\Omega)}.$
Then by (4,5), we have
$\|F_{h}(\chi_{h})\|_{L^{p}(\Omega)}\leq C\big(\|\overline{\nabla}_{h}(\nabla\chi_{h})-D^{2}u\|_{L^{p}(\Omega)}+\|\chi_{h}-u\|_{W^{1,p}(\Omega)}\big)$
$\leq$ $C\big(\|\overline{\nabla}_{h}(\nabla(\chi_{h}-u))\|_{L^{p}(\Omega)}+\|\overline{\nabla}_{h}(\nabla u)-\nabla(\nabla u)\|_{L^{p}(\Omega)}+\|\chi_{h}-u\|_{W^{1,p}(\Omega)}\big).$
By Definition 2.1 of the discrete gradient operator $\overline{\nabla}_{h}$, we have
$\|\overline{\nabla}_{h}(\nabla\chi_{h})-D^{2}u\|_{L^{p}(\Omega)}\leq C\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)},$
$\|\overline{\nabla}_{h}(\nabla u)-\nabla(\nabla u)\|_{L^{p}(\Omega)}=\|\overline{P}_{h}D^{2}u-D^{2}u\|_{L^{p}(\Omega)}.$
Therefore, (90) holds.
We can conclude that proof is complete. ∎
Acknowledgements The author would like to thank Professor Rob Stevenson for the fruitful discussion of this paper between them when he visited University of Amsterdam in 2025. The author was supported by Liu Bie Ju Centre for Mathematical Sciences (project no. 9600007, 9360109, 9360020).